### v1

# Capturing Stable HDR Videos Using a Dual-Camera System

Abstract
In HDR video reconstruction, exposure fluctuations in reference images from alternating exposure methods often result in flickering. To address this issue, we propose a dual-camera system (DCS) for HDR video acquisition, where one camera is assigned to capture consistent reference sequences, while the other is assigned to capture non-reference sequences for information supplementation. To tackle the challenges posed by video data, we introduce an exposure-adaptive fusion network (EAFNet) to achieve more robust results. EAFNet introduced a pre-alignment subnetwork to explore the influence of exposure, selectively emphasizing the valuable features across different exposure levels. Then, the enhanced features are fused by the asymmetric cross-feature fusion subnetwork, which explores reference-dominated attention maps to improve image fusion by aligning cross-scale features and performing cross-feature fusion. Finally, the reconstruction subnetwork adopts a DWT-based multiscale architecture to reduce ghosting artifacts and refine features at different resolutions. Extensive experimental evaluations demonstrate that the proposed method achieves state-of-the-art performance on different datasets, validating the great potential of the DCS in HDR video reconstruction. The codes and data captured by DCS will be available at https://github.com/zqqqyu/DCS.
Index Terms:
HDR video, dual-camera system, multi-exposure image fusion
I Introduction
Capturing the full range of illumination in high dynamic range (HDR) scenarios is a challenging task for standard digital cameras [1, 2]. Thanks to the multi-exposure fusion (MEF) technology [3, 4, 5], we can successfully capture HDR images in static scenarios by fusing multiple images of different exposure times. However, capturing HDR videos with a standard digital camera in dynamic scenes has not yet been effectively resolved. Early studies concentrated on designing hardware systems for directly capturing HDR video, including internal/external beam splitters [6, 7, 8], specialized sensors [9, 10], modulo cameras [11] and neuromorphic cameras [12], etc. Although these solutions show good potential for capturing high-quality HDR videos, their high cost and bulky design severely limit their widespread adoption.
Figure 1: Qualitative comparison between SAFNet and AE methods. Existing AE methods result in flickering artifacts (luminance fluctuations exceeding 30). In contrast, our system ensures consistent illumination across reference frames, enabling temporally consistent HDR reconstruction results (luminance fluctuations limited to under 1).
To address the limitations of these issues, researchers [13, 14, 15] have turned to computational methods that leverage inexpensive standard digital cameras, providing a promising direction for HDR video reconstruction. With the rapid advancement of deep learning, researchers make efforts to solve HDR imaging problems in dynamic scenes with deep neural networks. These methods can be broadly categorized into two types. One approach follows the MEF attempts, aiming to reconstruct an HDR image by fusing misaligned multi-exposure images. They typically use a specific exposure image as a reference frame for HDR reconstruction. As a result, capturing HDR video using these methods often requires a large number of additional exposure images, making it difficult to meet the real-time requirements of video capture. To overcome this limitation, alternating exposure (AE) methods [16, 17, 18, 19, 20] have been proposed, which allow varying exposures as reference frames and eliminate the need for additional exposure images. However, aligning the brightness of reference frames of different exposures and achieving consistent and stable lighting in the reconstruction video is still challenging. Although existing AE methods have achieved excellent performance on public benchmarks, they struggle to handle complex real-world scenes and lighting variations (shown in Fig. 1).
Generally, both existing MEF methods and AE methods struggle to achieve high-quality video capture in dynamic scenes. Compared to AE methods, MEF methods could produce HDR image sequences with stable and consistent lighting because of their use of a fixed-exposure reference. Noticing the wide use of multi-camera system in consumer electronic devices, it is possible to introduce multi-camera system for high-quality HDR video capture. Based on these insights, we develop a dual-camera system (DCS) for capturing high-quality HDR videos. In the DCS, one camera is assigned to capture reference images, while the other is assigned to capture non-reference images for MEF. In this way, the DCS could effectively mitigate the temporal inconsistencies associated with AE methods while avoiding the acquisition latency inherent in MEF.
Within DCS, HDR video reconstruction with MEF faces greater challenges than static images for two main reasons. First, exposures of input frames become uncertain. Second, the backgrounds from non-reference images are no longer aligned with the reference frame. Existing MEF methods can be categorized into two main approaches: optical-flow-based and attention-based. The former relies on optical flow to align inputs before fusion, which often suffers from a low accuracy in estimating optical flow. On the other hand, attention-based methods, while offering greater flexibility, mainly depend on the quality of the attention maps for fusion. Therefore, existing attention-based MEF methods encounter two main issues in reconstructing HDR videos within DCS. First, they fail to adequately leverage exposure information, relying primarily on texture and luminance to calculate attention maps, which may suffer from the various exposure settings. Second, they hardly consider the importance of reference frames, particularly in the computation of attention maps, where equal weights are often assigned to both reference and non-reference frames. This leads to attention being skewed by non-reference inputs, especially in complex, dynamic scenes, thereby hindering the effective elimination of artifacts.
Against the above limitations, we propose an exposure-adaptive fusion network (EAFNet) to achieve robust and generalized HDR reconstruction. Our EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a reconstruction subnetwork. The pre-alignment subnetwork consists of a global luminance alignment (GLA) and an exposure-guided feature selection module (EFSM) to leverage exposure information, enhancing meaningful details and ensuring that the most relevant features are preserved despite varying exposure levels. The asymmetric cross-feature fusion subnetwork aims to explore reference-dominated attention maps, align cross-scale features, and integrate cross-feature information to improve the fusion. Additionally, a multiscale reconstruction subnetwork is introduced to suppress ghosting artifacts and enhance fine details. Summary, our contributions can be summarized as follows:
  • We propose a new HDR video generation paradigm that uses a fixed medium-exposure sequence as the temporal reference, while leveraging a second varying-exposure stream to enrich dynamic range. This design fundamentally improves temporal consistency and reconstruction stability.
  • We designed and implemented an asynchronous dual-camera system to validate the feasibility of our proposed paradigm and bridge the gap between algorithmic design and practical deployment. Furthermore, existing MEF methods can be seamlessly applied within our system, achieving temporally consistent reconstruction.
  • Built upon MEF methods, we propose a novel model design, EAFNet. The proposed pre-alignment subnetwork leverages exposure information and explores the intrinsic properties of the images, ensuring that the most relevant features receive attention despite varying exposure levels. An asymmetric cross-feature fusion subnetwork is introduced to explore reference-dominated attention maps, achieving smoother and more consistent fusion through cross-scale alignment and cross-feature interaction.
  • Our EAFNet achieves state-of-the-art results on public benchmarks. Both qualitative and quantitative experiments demonstrate the effectiveness of our EAFNet and DCS.
II RELATED WORK
II-A HDR Image Reconstruction.
II-A1 Traditional Methods.
Traditional multi-exposure image fusion HDR imaging methods usually solve the ghosting problem in the image domain, and currently there are two main types of solutions:
Motion-aligned-based. These methods focus on a global alignment of the LDR input and then reject the unaligned pixels before image fusion to achieve de-ghosting. Bogoni et al. [4] used optical flow to estimate motion vectors between LDR images and warp the moving pixels accordingly. Kang et al. [16] improved the alignment results by using the exposure time information to transform the intensity of LDR images to the brightness domain before aligning them with the optical flow method. Sen et al. [5] proposed a patch-based energy minimization method that integrates alignment and HDR reconstruction into a joint optimization. Hu et al. [21] optimized alignment using luminance and gradient consistency in the transform domain. Hafer et al. [22] proposed an energy minimization approach to jointly compute HDR radiance and displacement fields. These motion-alignment methods heavily rely on the accuracy of the alignment algorithms.
Motion-rejection-based. Motion-rejection methods detect and exclude motion regions before fusion. Grosch et al. [23] identified motion regions based on intensity differences, while Jacobs et al. [13] used weighted variance to detect ghosting. He et al. [24] refined motion detection by combining joint probability density with graph-cut-based energy minimization. Zhang et al. [25] removed motion pixels using image gradient information, while Lee et al. [26] and Oh et al. [27] introduced rank minimization to detect motion and reconstruct HDR images. However, discarding motion regions often compromises information recovery in static areas.
II-A2 Deep Learning Approaches.
With the rapid advancement of deep learning, neural network-based methods for HDR image reconstruction have gained significant attention and have emerged as the prevailing approach in the field, these can be categorized primarily into the following types:
Flow-based Alignment. Kalantari et al. [28] first applied CNNs to multi-exposure HDR imaging, using optical flow to register non-reference LDR images to a reference image. Since LDR images contain not only motion regions but also saturated regions, which do not satisfy the assumption of luminance consistency of the optical flow method, Peng et al. [29] and Prabhakar et al. [30] suggested that the typical optical flow algorithm may lead to large misalignment errors. Therefore, they used optical flow networks (e.g., FlowNet) to achieve alignment. SAFNet [31] refined valuable regions and estimated motion to produce high-quality HDR images. However, when long-distance motions occur in the LDR images, these methods still struggle to estimate the accurate flow for alignment.
Direct Feature Concatenation. Wu et al. [32] introduced a flow-free approach, treating HDR imaging as an image transformation problem. They first applied homography to align multiple frames globally and then used a multiscale network to map LDR images to HDR images directly. A similar approach was also employed in [33]. Niu et al. [34] and Li et al. [35] used Generative Adversarial Networks (GANs) for HDR reconstruction from multi-exposure LDR images with significant motion. However, they still face challenges in handling extreme motion and dynamic scenes.
Attention-based Alignment. AHDRNet [36] introduced the attention mechanism into multi-exposure HDR reconstruction for dynamic scenes, coupling it with a densely connected dilated residual module to achieve alignment and fusion. Liu et al. [37] introduced a Pyramid Cascading Deformable (PCD) module, using multiscale deformable convolutions for feature alignment in LDR images. Chen et al. [38] proposed novel attention mechanisms and progressive texture fusion to improve reconstruction. DomainPlus [39] utilized multiscale attention with fronted fusion to perform coarse-to-fine fusion. HDR-Trans et al. [40] added Transformer to the CNN-based HDR architecture, combining the advantages of CNN in extracting local information and Transformer in capturing global information. Chen et al. [41] proposed a transformer model for HDR imaging with three steps: alignment, fusion, and reconstruction. Yan et al. [42] proposed a HyHDRNet consisting of a content alignment subnetwork and a transformer-based fusion subnetwork to improve performance. SCTNet [43] integrated spatial and channel attention in a transformer-based network to improve semantic coherence. Despite advances in these approaches, they still generate undesirable ghosts in high-contrast regions.
II-A3 Dual-lens-based system
Dual-lens HDR imaging leverages paired images with different exposures to synthesize high-quality HDR content. Dong et al. [44] employ a soft-warping step to align long-exposure features to the short-exposure reference, Li et al. [45] propose a dual-stage model: a guided 3D exposure correction network estimates a coarse HDR image from the short exposure input, followed by a guided denoising transformer that refines the result using long-exposure features. While effective under controlled conditions, both frameworks assume the only misalignment present is 1D parallax between image pairs. This assumption is embedded in their data collection and rectification procedures, simplifying alignment to left–right disparities but limiting their ability to model more complex distortions. In dynamic scenes, exposure time mismatch introduces additional misalignments: high-exposure frames exhibit motion blur not present in low-exposure frames.
II-B HDR Video Reconstruction
Dedicated Hardware Solutions. Dedicated hardware solutions, such as single-pixel imaging [46], scanline exposure/ISO [9], internal [8] or external beam splitters [7], as well as residue cameras [11] and neuromorphic vision cameras [12], are capable of rapidly and efficiently generating detailed HDR images or videos on specialized devices. However, the complexity and high cost of these hardware solutions limit their widespread adoption.
Traditional Methods. Similar to multi-exposure HDR image fusion, most methods utilize alternating exposure sequences to align multiple frames for HDR video reconstruction. Kang et al. [16] introduced the first algorithm in this category, employing global and local registration to align multi-exposure images at each time step and merge them into a single HDR frame. Kalantari et al. [17] employed patch-based optimization to reconstruct missing images from different exposures. Mangiat et al. [17] applied block-based motion estimation with refined motion vectors and later improved this method with HDR filtering to eliminate block artifacts [47]. Additionally, Gryaditskaya et al. [18] proposed an adaptive weighting algorithm to minimize motion artifacts. Li et al. [48] proposed a statistical method to avoid exact alignment estimation. However, these methods are typically time-consuming and often introduce visible artifacts.
Deep Learning Methods. Kalantari et al. [49] proposed the first deep learning-based method for generating HDR video from alternating exposure sequences. Chen et al. [19] developed a two-stage coarse-to-fine framework that integrated deformable convolutions with HDR video reconstruction. LAN-HDR [50] aligned adjacent frames with the reference frame using luminance-based attention scores. HDRFlow [51] designed an efficient flow network with multiscale large kernels to handle large motions. Cui et al. [20] introduced exposure completion for neural HDR rendering, using feature interpolation and a novel decoding architecture to improve HDR video reconstruction. Nonetheless, the varying exposures of the reference frames create a significant challenge in maintaining temporal consistency in video reconstruction.
Figure 2: Visualization of our dual-camera system. The primary camera captures continuous medium-exposure sequences as reference for temporal consistency, while the secondary camera alternates between low- and high-exposure to provide complementary information for reconstruction.
III DUAL-CAMERA SYSTEM FOR HDR VIDEO
Overview. A mainstream solution for HDR video reconstruction is generating video from image sequences captured at alternating exposure levels. However, this approach inherits the exposure limitations of the reference frame, often resulting in visible artifacts and temporal flickering in the generated video, significantly constraining its practical application. To address this, we propose a dual-camera system (As shown in Fig. 2 (b)), where the primary camera continuously captures medium-exposure images to ensure temporal consistency, while the secondary camera captures low-exposure and high-exposure images to provide complementary information.
III-A Dual-Camera System
System Configuration. Our system employs two MV-CS032-10GC industrial cameras, which provide low-level access to operational controls for precise adjustment of exposure time and shooting strategies. Both cameras are equipped with identical lenses and are mounted side by side, secured by a mechanical support structure to ensure their relative positions remain stable throughout the imaging process.
Dual-Camera Calibration. To eliminate distortion and ensure alignment, we perform camera calibration using a checkerboard pattern. Intrinsic and extrinsic parameters are computed from images captured by the dual cameras using the camera calibration toolbox [52]. The calibration parameters are then applied to register the images, resulting in the final aligned images. This reduces potential errors and improves the subsequent processing steps.
Video Capture. Both cameras are triggered to start capturing images simultaneously, but subsequent image capture can occur asynchronously. One camera captures the medium-exposure reference sequence $L_{reference}(t)$, while the other captures the alternating exposure non-reference sequence $L_{non-reference}(t)$ with low and high exposure. Both cameras are set to the same resolution of $w\times h$. Due to different exposure time settings, the videos captured by the two cameras are not timestamp-synchronized. However, our method does not require precise synchronization. We use timestamp information to roughly group the data, where each group consists of frames from $L_{reference}(t)$ and two high- and low-exposure images from $L_{non-reference}(t)$ (e.g., with exposure times of (-2, 0, 2 EV), every four frames from the reference sequence, along with one pair of high- and low-exposure images, form a single group). Each frame, along with the two high- and low-exposure images, is used as input to the network. The final reconstructed HDR video matches the frame rate of $L_{reference}(t)$. To ensure a fair comparison, we capture test datasets for both AE and MEF using our system. Detailed information can be found in Sec. V-B, and the datasets will be available.
Essentially, our system can be viewed as a multi-exposure HDR fusion task for dynamic scenes, where luminance alignment between neighboring frames is unnecessary due to consistently exposed reference frames. Existing MEF methods can be seamlessly extended to achieve consistent exposure results. To address challenges from video data and the variability of real-world scenes, we propose SAFNet, a more robust network, detailed in Section IV.
Figure 3: The architecture of EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a restoration subnetwork. We introduce GLA and EFSM to leverage exposure information, explore the intrinsic properties of the images, and help preserve finer details across varying exposures. The asymmetric cross-feature fusion subnetwork improves image fusion by aligning cross-scale features and performing cross-feature fusion. The reconstruction subnetwork adopts a DWT-based multiscale architecture to reduce ghosting and refine features at different resolutions.
III-B Discussion.
The proposed DCS offers several advantages over traditional single-camera AE methods for HDR video reconstruction:
Stable Lighting Conditions. The DCS provides more stable exposure across frames, avoiding the fluctuations typical in alternating exposure systems.
Potential in Low-Light Conditions. The DCS performs better in low-light environments, where long exposure times in a single-camera setup can lead to artifacts and delays. The dual-camera approach mitigates these issues by distributing exposure tasks between the cameras.
Offline HDR Video Reconstruction. The DCS setup supports offline HDR video reconstruction, allowing continuous high-quality video capture without real-time processing, ensuring high-quality capture without real-time processing delays. In contrast, AE methods require real-time reconstruction to avoid flickering, which can introduce undesirable artifacts if not processed immediately.
Independent Automatic Exposure. Each camera in the DCS can independently apply established automatic exposure strategies, eliminating the need for recalibration with each frame and simplifying real-time operation. In contrast, AE methods require constant adjustments to exposure settings, complicating real-time capture.
Seamless Integration with Encoding Strategies. The dual-camera framework integrates easily with existing video encoding and decoding strategies. The secondary camera can capture frames at a low frame rate (e.g., 1fps or lower), minimizing additional bandwidth requirements, thus ensuring efficient video transmission.
IV METHOD
IV-A Overview
Given a set of LDR images of a dynamic scene $\left\{L_{1},L_{2},\ldots,L_{n}\right\}$ with different exposures, HDR imaging aims to generate an HDR image $H$ aligned to a prescribed reference image. Fig. 3 shows an overview of the proposed EAFNet. Our proposed EAFNet consists of three main subnetworks: the pre-alignment subnetwork, the asymmetric cross-feature fusion subnetwork, and the restoration subnetwork.
IV-B Pre-alignment
The pre-alignment subnetwork aims to extract meaningful details from the non-reference image by leveraging exposure information. Global luminance alignment (GLA) is introduced to maintain consistent luminance across varying exposures, while the exposure-guide feature selection module (EFSM) explores the correlation between inputs and exposure to focus on the most relevant regions for fusion.
Global Luminance Alignment. Following [53, 36], we employ three LDR images $\left\{L_{l},L_{m},L_{h}\right\}$ (sorted by their exposure length) as input. The images first undergo gamma correction [53], defined as:
$H_{i}=L_{i}^{\gamma}/{t_{i}},\ \ i=l,m,h$ (1)
where $\gamma>1$ denotes the gamma correction parameter and $t_{i}$ denotes the exposure time. The indices $l,m,h$ correspond to low, mid, and high exposure settings, respectively, with $m$ serving as the reference image. In this work, the default gamma parameter $\gamma$ is set to 2.2. Then, we will get a corresponding set of $\left\{H_{i}\right\}$.
Due to varying luminance levels at different exposure times, captured images may exhibit color rendering deviations from real-world colors, with low-exposure images losing shadow details and high-exposure images losing highlight details. Therefore, variations in exposure times certainly affect the overall color accuracy of the input LDR images. To address this, we propose a GLA in the pre-alignment subnetwork. GLA aims to correct the luminance of LDR images in the sRGB domain, producing consistent brightness features across varying exposures and enhancing detail preservation. Specifically, the LDR images are mapped to the luminance range of the reference image and achieve global luminance alignment:
$G_{i}=clip(L_{i}*(mean(L_{i})/mean(L_{m})),\ i=l,h)$ (2)
where $G_{i}$ denotes the corresponding luminance-aligned output, $mean(\cdot)$ denotes the mean luminance value, and $clip$ is a function that keeps the output in the range [0, 1]. Notably, since $L_{m}$ is the reference image, $G_{m}=L_{m}$.
Figure 4: The pre-alignment subnetwork is composed of two parts: global luminance alignment (GLA) and exposure-guided feature selection module (EFSM). The input is divided into multiple scales using a $3\times 3$ convolution, with a shared-weight feature selection block applied at each scale.
As shown in Fig. 4, given a set of LDR images with varying exposure times, the non-reference images can be aligned to the reference image using GLA. Finally, we can obtain 3-pairs of 9-channel tensor $L_{in}=\left(\{L_{l},H_{l},G_{l}\},\{L_{m},H_{m},G_{m}\},\{L_{h},H_{h},G_{h}\}\right)$ as the input to feed into the EFSM module.
Exposure-Guided Feature Selection Module. Existing methods primarily focus on texture for fusion, overlooking the crucial role of exposure time information, which may result in suboptimal fusion and color inconsistencies. To overcome this limitation, we propose an EFSM to explore exposure information, enhancing meaningful details and ensuring that the most relevant features are preserved despite varying exposure levels.
For a set of LDR images with corresponding exposure times of $\left\{{{t_{l}},{t_{m}},{t_{h}}}\right\}$, we standardize and relativize the input exposure times using the following formulation:
${e_{{}_{l}}}={\log_{2}}\left({{{t_{l}}}/{{t_{m}}}}\right)\cdot c,\ \ \ {e_{{}_{h}}}={\log_{2}}\left({{{t_{h}}}/{{t_{m}}}}\right)\cdot c$ (3)
where ${e_{{}_{l}}}$ and ${e_{{}_{h}}}$ denote the relative exposure inputs of the low exposure branch and high exposure branch, respectively. $c$ is the hyper-parameter set to 0.1.
As shown in Fig. 4, we adopt a multiscale architecture to progressively refine feature representations across different exposure levels. The framework initiates with multi-level feature extraction: given input $L_{in}$, three parallel $3\times 3$ convolutional layers generate initial feature maps $\{f_{l},f_{m},f_{h}\}$ corresponding to low-, mid-, and high-exposure features, respectively. For subsequent processing clarity, we formally define $f^{nr}\triangleq\{f_{l},f_{h}\}$ as non-reference features and $f^{r}\triangleq f_{m}$ as reference features. Assuming there are $N$ scales in total, the reference and non-reference features of the $i$-th scales are denoted as $f_{i}^{r}$ and $f_{i}^{nr}$, respectively ($i=1$ denotes the original scale), $f_{i}^{nr}$ are fed to the shared weights exposure-guided feature selection block (EFSB) and $f_{i}^{r}$ are fed to the shared weights feature selection block (FSB). Specifically, the features undergo global average pooling, followed by two fully connected layers and a sigmoid activation to produce the feature-guide modulation coefficients $v^{nr}\in\mathbb{R}^{1\times 1\times 64}$ and $v^{r}\in\mathbb{R}^{1\times 1\times 64}$. Simultaneously, relative exposure inputs $e$ are each passed through fully connected layers, resulting in the exposure-guide modulation coefficients $v^{t}\in\mathbb{R}^{1\times 1\times 128}$.
To dynamically select and refine features from multiple input channels, the exposure-guide modulation coefficients $u^{nr}\in\mathbb{R}^{1\times 1\times 64}$ are computed as follows:
$u^{nr}=\sigma\left(FC_{3}\left(<v^{nr},v^{t}>\right)\right)$ (4)
where <,><,>< , > denotes concatenation, $\sigma$ represents sigmoid activation, $FC_{3}$ denotes three fully connected layers. The final output of EFSM can be obtained as:
$\hat{f_{i}^{r}}=f_{i}^{r}\cdot v^{r},\ \ \ \ \ \ \hat{f_{i}^{nr}}=f_{i}^{nr}\cdot u^{nr}$ (5)
IV-C Asymmetric Cross-feature Fusion
While the non-reference image may provide more details, it also poses a higher risk of introducing ghosting artifacts, especially in open scenes. To address this, we propose an asymmetric cross-attention block, which enhances the contribution of the reference input during attention computation and mitigates the potential for ghosting artifacts.
Figure 5: The structure of asymmetric cross-feature block. It employs asymmetric cross-attention to align reference and non-reference features, integrating coarse-to-fine guidance for enhanced reference-dominated feature fusion and improved HDR reconstruction.
Generally, we propose a multi-scale strategy to calculate the attention in a coarse-to-fine manner. At each scale, asymmetric cross-attention (ACA) is utilized to compute local attention. Specifically, as shown in Fig. 5, the proposed ACA first utilizes the non-reference feature as an agent, facilitating the exchange of information between the patch tokens of the reference feature and then back project it to its non-reference branch. To prioritize mid-exposure image features and encourage more effective exploration of feature interactions across different exposures, we introduce the self-information during the exchange process. Following [54], the attention matrix is thus computed as follows:
$Q_{i}=\phi(\hat{f}_{i}^{r})W_{i}^{q},\ K_{i}=\phi(\hat{f}_{i}^{nr})W_{i}^{k},\ V_{i}=\phi(\hat{f}_{i}^{nr})W_{i}^{v}$ (6)
$A_{i}=softmax((Q_{i}+K_{i})K_{i}^{T}/\sqrt{C})$ (7)
where $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{C\times C}$ are learnable weight matrices of the $i$-th scales, $\phi(\cdot)\in\mathbb{R}^{N\times C}$ denotes the unfolding operation ($N$ refers to the number of patches). Then, the aligned non-reference features $f^{anr}_{i}$ are obtained by applying a folding operation to the result of the multiplication of $A_{i}$ and $V_{i}$.
To leverage attention from coarser scales and local details at the current scale, we introduce the cross-scale guidance feature $f^{guid}$ to achieve coarse-to-fine feature fusion and interaction:
[…]
where $f^{guid}_{i}$ represent guidance features at $i$-th scales, $U$ represents pixel shuffle upsampling, $C_{v}$ denotes a $3\times 3$ convolution. As shown in Fig. 5, $f^{guid}_{i}$ learn reference-dominated attention from the previous scale, facilitating the alignment of features across scales. The output of asymmetric cross-attention block can be expressed as:
$\tilde{f}^{nr}=C_{v}(<f^{anr}_{1},U(f^{anr}_{2}),f^{guid}_{1}>)$ (9)
Finally, we generate a spatial attention map with $\tilde{f}^{nr}$ and $f^{r}$ and get the final output:
$f^{out}=<\sigma(C_{v}(Relu\left(C_{v}\left(<f^{r},\tilde{f}^{nr}>\right)\right)))\cdot\tilde{f}^{nr},f_{m}>$ (10)
IV-D Restoration
The restoration subnetwork adopts a frequency-based approach to effectively mitigate ghosting artifacts in HDR images. As shown in Fig. 3, the subnetwork is designed with a multiscale architecture that progressively refines features at different resolutions. The architecture particularly emphasizes feature correction at lower scales, as these scales capture finer high-frequency details, such as textures and small features, which are essential for addressing the residual ghosting artifacts.
At the lower scales, the features undergo Discrete Wavelet Transform (DWT) to decompose the image into multi-resolution subbands. Following this, multiple Learnable Bandpass Filter (LBF) modules [55, 39] are applied to correct each frequency component. The output is then passed through the Inverse Wavelet Transform (IWT) to reconstruct the feature maps $f_{\times 1}$ in the feature domain. At higher scales, after applying DWT, a $1\times 1$ convolution is used instead of multiple LBF modules. The feature maps at higher scales primarily contain global structural information, reducing the need for complex high-frequency corrections.
Finally, we combine the high-frequency details from lower scales with the global information from higher scales, leading to improved HDR image restoration with reduced ghosting artifacts:
$H_{out}=\sigma(Conv(U(U(f_{\times 4})+f_{\times 2})+f_{\times 1}))$ (11)
where $U$ denotes upsampling, $f_{\times 4},f_{\times 2},f_{\times 1}$ are the output features at different scales, $H_{out}$ represents the final HDR image output.
IV-E Loss Function
HDR images are typically displayed after tone mapping. To enhance training performance, we apply the $\mu$-law transformation, as proposed in [28], to map the HDR image from the linear domain to the tone-mapped domain, thereby improving model performance during training, the $\mu$-law transformation can be expressed as:
$\tau(H)=\frac{\ln(1+\mu H)}{\ln(1+\mu)}$ (12)
where $\mu$ is a hyperparameter set to 5000.
We adopt the L1 loss as the basic loss function for pixel-wise supervision. To correct each frequency component and enhance the training of the restoration subnetwork, we adopt the dilated advanced Sobel loss (D-ASL) loss function [56] as an additional function, which is defined as:
$\mathcal{A}\mathcal{S}\mathcal{L}(\widehat{Z},Z)=\frac{1}{N}\sum\mid\operatorname{Sobel}^{\star}(\widehat{Z})-\text{ Sobel }^{\star}(Z)\mid\$ (13)
${\cal D}-{\cal A}{\cal S}{{\cal L}^{\left\{{{d_{1}},{d_{2}},\cdots,{d_{n}}}\right\}}}=\sum\limits_{i=1}^{n}{{\cal A}{\cal S}{\cal L}\left|{{}_{dilation\_rate={d_{i}}}}\right.}$ (14)
where the $N$ is the batchsize of training, $Z$ and $\widehat{Z}$ represent the groundtruth and the output image predicted by network after the $\mu$-law function, $\operatorname{Sobel}^{\star}$denotes the advanced sobel filtering, and ${{\cal A}{\cal S}{\cal L}\left|{{}_{dilation\_rate={d_{i}}}}\right.}$ denotes the ASL with a dilation rate of $d_{i}$. Following [56], we apply the reference settings that $D=\left\{{1,2,3}\right\}$ to formulate the D-ASL. Then the final loss function can be expressed as:
$\operatorname{Loss}(\widehat{Z},Z)=\mathcal{L}1(\widehat{Z},Z)+\lambda\cdot\mathcal{D}-\mathcal{AS}\mathcal{L}^{D}(\widehat{Z},Z)$ (15)
where $\lambda$ = 0.25 is a hyper-parameter to balance the L1 loss and D-ASL.
V EXPERIMENTS
V-1 Settings
The proposed method is implemented in PyTorch on an NVIDIA RTX4090 GPU. For the convolutional layers, we employ 64 kernels of size 3$\times$3 with a stride of one, applying zero padding to preserve the dimensions of the resulting feature maps. The optimization is performed using the Adam optimizer [57] with an initial learning rate of $10^{-4}$, and the training concludes when the learning rate reaches $10^{-6}$. To train the model, both the LDR and corresponding images are cropped into 256 $\times$ 256 patches. A batch size of 16 is utilized during training. Patches are randomly rotated for data augmentation to avoid overfitting during the training stage.
V-2 Datasets
We employ four benchmark datasets for evaluation: Prabhakar’s dataset [30], Kalantari’s dataset [28], Tursun’s dataset [58] and Sen’s dataset [5]. Details of these datasets are described as follows:
1). Prabhakar’s Dataset. Including 582 varying exposure images with corresponding deghosted HDR images, with exposure values between -3EV to +3EV.
2). Kalantari’s dataset. Including 74 training samples and 15 testing samples. Each consists of three LDR images, and the exposure values belong to {-2, 0, +2} or {-3, 0, +3}.
3). Tursun’s dataset and Sen’s dataset. Lacking ground truth, is used to assess the generalization ability of the models.
V-3 Metrics
We quantitatively evaluate the predicted HDR images by measuring PSNR and SSIM [59] in both the linear domain (-L) and the tone-mapped domain (-$\mu$) [28]. The tone mapping operation is described in Sec. IV-E. Additionally, we compare the results using HDR-VDP-2 [60] for conforming the human visual perception. We also quantitatively evaluated the luminance consistency of the video by testing the Luminance Standard Deviation (LSD), temporal SSIM (t-SSIM) and average luminance ($L_{avg}$):
  • Average Luminance ($L_{avg}$):
    $L_{avg}=\frac{1}{N}\sum_{i=1}^{N}L_{i}$ (16)
    where $N$ represents the number of the frame, $L_{i}$ represents the mean luminance value of frame $i$.
  • Luminance Standard Deviation (LSD):
    $\text{LSD}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(L_{i}-L_{avg})^{2}}$ (17)
    where $L_{i}$ is the mean luminance of frame $i$.
V-A Comparison with state-of-the-art methods
We quantitatively and qualitatively compare our EAFNet to several state-of-the-art MEF methods including Kalantari [49], AHDR [36], Prabhakar [61], HDR-Transformer [40], SCTNet [43], DomainPlus [39], SAFNet [31], and several state-of-the-art AE methods including HDRFlow [51], LAN-HDR [50], and DeepHDRVideo [19].
TABLE I: Comparison results on Kalantari’s dataset and Prabhakar’s dataset. The best results are highlighted and the second best results are underlined.
Models 17’CGF 19’CVPR 20’ECCV 22’ECCV 22’MM 23’ICCV 24’ECCV
Kalantari AHDRNet Prabhakar HDR-Trans DomainPlus SCTNet SAFNet Ours
train and test on Kalantari’s dataset
PSNR-µ 42.74 43.77 43.08 44.28 44.02 44.13 44.61 44.69
PSNR-L 41.22 41.35 41.68 42.88 41.28 42.12 43.09 42.19
SSIM-µ 0.9877 0.9907 / 0.9916 0.9910 0.9916 0.9918 0.9920
SSIM-L 0.9848 0.9859 / 0.9884 0.9864 0.9890 0.9892 0.9895
HDR-VDP-2 60.51 62.30 62.21 66.03 62.91 66.65 66.93 68.35
train and test on Prabhakar’s dataset
PSNR-µ 35.63 38.61 38.30 41.31 40.38 41.23 40.18 41.80
PSNR-L 32.50 35.26 34.98 39.44 38.08 38.75 37.90 40.13
SSIM-µ 0.9613 0.9663 0.9702 0.9726 0.9698 0.9724 0.9705 0.9731
SSIM-L 0.9692 0.9794 0.9781 0.9885 0.9872 0.9881 0.9865 0.9895
HDR-VDP-2 59.42 61.14 / 63.01 62.12 62.29 62.04 63.53
TABLE II: Cross-dataset evaluation on Kalantari’s dataset and Prabhakar’s dataset. The best results are highlighted and the second best results are underlined.
2019 2022 2022 2023 2024
Models AHDRNet HDR-Trans DomainPlus SCTNet SAFNet Ours
train on Kalantari’s dataset, test on Prabhakar’s dataset
PSNR-µ 33.96 34.07 32.64 33.83 38.00 39.26
PSNR-L 32.46 36.62 30.4233 30.95 34.65 35.99
SSIM-µ 0.9601 0.9675 0.9046 0.9584 0.9597 0.9707
SSIM-L 0.9542 0.9656 0.9094 0.9521 0.9793 0.9848
train on Prabhakar’s dataset, test on Kalantari’s dataset
PSNR-µ 40.03 41.38 41.15 40.88 40.86 42.02
PSNR-L 36.71 39.21 38.18 37.59 37.50 39.38
SSIM-µ 0.9855 0.9890 0.9873 0.9892 0.9882 0.9903
SSIM-L 0.9758 0.9873 0.9837 0.9842 0.9810 0.9870
Figure 6: Visual comparison on Kalantari’s dataset. While other methods suffer from detail loss, overexposed patches, and color inconsistencies, our approach successfully reconstructs overexposed details and preserves natural, consistent color.
Figure 7: Visual comparison on Prabhakar’s dataset. Most existing methods suffer from ghosting artifacts, our EAFNet mitigates these issues.
Experiments on Prabhakar’s datasets and Kalantari’s datasets.
Intra-Validation. Our DCS can be viewed as a MEF task, so we conduct a comparison with the state-of-the-art MEF method using the publicly available MEF dataset. The quantitative results on both datasets are shown in Table I, our EAFNet achieves superior performance across all datasets. Specifically, on the Kalantari dataset, our method surpasses the second-best approach in terms of PSNR-$\mu$ by 0.08 dB, while on the Prabhakar dataset, it achieves a more substantial improvement of 0.49 dB. The visual comparisons in Fig.  7 illustrate that most existing methods still suffer from noticeable ghosting artifacts. In contrast, the proposed EAFNet effectively mitigates these issues. Moreover, in overexposed regions (Figure 6), we find that the Kalantari method leads to detail loss, while AHDRNet and HDR-Trans exhibit large white overexposed patches in wall areas. Additionally, methods such as DomainPlus still struggle with color inconsistencies. In contrast, our approach successfully reconstructs most of the overexposed details while preserving natural and consistent color reproduction.
Figure 8: Visual comparison on Tursun’s dataset. Existing methods often produce artifacts due to luminance contrast between reference and non-reference images, while SAFNet effectively suppresses these artifacts.
Figure 9: Visual comparison on Sen’s dataset. Existing methods suffer from misalignment or detail loss when the reference image is over-exposed, while our method reconstructs clearer and more accurate textures.
Cross-Validation. The cross-testing on the two datasets serves to mitigate any potential bias and offers a comprehensive understanding of the model’s reliability across varying conditions. As shown in Table II, we achieve superior performance in cross-dataset evaluation, demonstrating the robustness and generalizability of our model.
Unsupervised-dataset. To assess the generalization performance of our SAFNet, we conduct qualitative comparisons on the Sen’s and Tursun’s datasets. As shown in Fig. 8, when there exists a significant luminance contrast between the reference and non-reference images, such as the case where white backgrounds in the reference image correspond to dark content in the non-reference frames, most existing methods tend to produce prominent visual artifacts. In contrast, the proposed SAFNet effectively suppresses artifacts, delivering perceptually plausible HDR reconstructions even in scenes with severe exposure disparity. In Fig. 9, When the reference image is over-exposed and misalignment exists between the reference and non-reference images, methods such as AHDRNet, DominPlus, and SAFNet exhibit incorrect alignment, while HDR-Trans and SCTNet suffer from varying degrees of detail loss. In contrast, our method is capable of reconstructing clearer and more accurate textures.
Figure 10: Visual comparison with MEF methods on self-captured videos. Our EAFNet reduces ghosting from motion (first row) and background misalignment (second row), recovering natural colors and fine details in dark regions (third row).
Figure 11: Visual comparison with AE methods on self-captured video dataset. Our dual-camera system reconstructs HDR video with consistent luminance.
TABLE III: Comparison on self-captured videos. * indicates MEF methods under the AE setup. The reported time refers to the inference time for a 128×128 patch on a single RTX 4090 GPU.
AE setup MEF setup with DCS
21’ICCV 23’ICCV 24’CVPR 23’ICCV 24’ECCV 22’ECCV 22’MM 23’ICCV 24’ECCV
Chen LAN-HDR HDRFlow SCTNet* SAFNet* HDRTrans DomainPlus SCTNet SAFNet Ours
LSD 0.1031 0.1003 0.0914 0.1027 0.0976 0.0093 0.0094 0.0093 0.0097 0.0093
t-SSIM 0.5913 0.6879 0.7352 0.6947 0.6862 0.8923 0.8991 0.8868 0.8851 0.9071
$L_{avg}$ 105.02 108.75 113.49 109.23 111.70 111.24 110.41 106.84 111.32 133.16
Runtime(ms) 22.23 21.69 2.6 11.17 3.2 21.6 7.44 11.17 3.2 4.3
V-B Experiments on self-captured videos
Overview. In this section, we explore the effectiveness of our DCS in real-world scenes and assess the capabilities of the proposed EAFNet in HDR video reconstruction. As detailed in Section III, real video data is captured using dual MV-CS032-10GC industrial cameras. Due to the alternating exposure method’s requirement for continuous alternating exposure frames, direct comparison with our approach under identical settings is not feasible. To ensure a fair comparison, we simultaneously capture two datasets with our system, ensuring minimal scene variation. The first camera captures alternating low, medium, and high exposure sequences, while the second camera continuously captures a medium exposure sequence. The AE method uses the alternating exposure sequence from the first camera as the test dataset, while our dual-camera framework takes low and high-exposure images from the first camera and a continuous medium-exposure sequence from the second. Both cameras have a resolution of $960\times 540$, and we select nine scenes spanning indoor, outdoor, day, and night conditions, each covering low-light, normal-light, and high-light scenarios with challenges such as high contrast, motion, and occlusion. The gain follows the scene variations, using exposure times of (-2, 0, 2 EV).
Compare with MEF methods. We compare our method with the state-of-the-art MEF methods including HDR-Transformer [40], SCTNet [43], DomainPlus [39], and SAFNet [31]. Firstly, we compare the performance of models trained on two different datasets, Prabhakar’s and Kalantari’s. As shown in Fig. 10 (third row), we find that models trained on Kalantari’s dataset exhibit unnatural textures in certain scenes. Therefore, all MEF results are tested using the model trained on Prabhakar’s dataset. During real video capture, complex lighting conditions, camera and object motion, and high-contrast scenes pose significant challenges for video reconstruction. As shown in the first row of Fig. 10, DomainPlus produces motion misalignment artifacts, while SAFNet and STCNet generate unnatural artifacts in some very dark regions. In contrast, our EAFNet effectively eliminates these artifacts, restoring more natural colors and brightness in darker regions, thus providing more detailed information. The second row of the figure demonstrates that for high-contrast scenes (where reference images are overexposed or areas with white backgrounds correspond to low/high exposure images of dark objects), existing methods fail to eliminate misalignment artifacts. However, our EAFNet can produce more natural reconstruction results, delivering a visually coherent output. The third row shows that our EAFNet provides more detail in the dark region, which leads to high luminance in Table III.
Compare with AE methods. Since the AE method and our method require training on different dataset settings, the AE method is trained on its public dataset following the AE training setup, while our method is trained on the publicly available Prabhakar’s dataset. The models are then tested on our dual-camera dataset. We compare the state-of-the-art AE methods including HDRFlow [51], LAN-HDR [50], and DeepHDRVideo [19]. As shown in Fig. 11 and Table III, for real-world test videos captured in our DCS, the luminance of the video reconstructed using the AE method exhibits abrupt changes, leading to noticeable flickering, which makes it impractical for real-world applications. In contrast, our DCS produces temporally consistent results, which suggests that the proposed system can effectively overcome the challenges of traditional AE methods.
Efficiency Analysis. The overall efficiency of HDR video processing is determined by both the acquisition throughput and the inference speed. Our method specifically targets the acquisition bottleneck by introducing an asynchronous dual-camera design, which decouples the capture of alternate exposure frames from the reference stream. This architecture effectively eliminates the latency caused by long high-exposure frames in traditional single-camera systems, enabling real-time HDR frame acquisition. Although our reconstruction model (EAFNet) is not specifically optimized for real-time inference, as shown in Table III, it still exhibits competitive runtime performance.
Dual-Camera Analysis. The use of a dual-camera system inevitably introduces parallax and temporal misalignment due to the baseline separation and potential desynchronization between the two cameras. In our prototype, the baseline is constrained by mechanical limitations but is minimized to the extent possible. Although some parallax is unavoidable, we find that most MEF-based methods in our system demonstrate reasonable robustness to mild parallax. The training data involves dynamic scenes with moving objects and occlusions, the model thus learns to handle motion-induced spatial misalignments, which share similar characteristics with parallax-related shifts in dual-camera setups. In challenging cases, such as those involving high local contrast, misalignment becomes more difficult to resolve. For instance, in the second row of Figure 10, the presence of strong contrast across exposure frames poses a severe test for existing methods. In typical dynamic scenes with such conditions, most existing approaches struggle to produce satisfactory results. In contrast, our proposed EAFNet effectively mitigates artifacts, preserving detail and structure even under these difficult situations. This robustness is achieved through a combination of a pre-alignment module and a reference-guided fusion strategy, which together reduce visible artifacts even in scenes with high contrast or occlusion.
In addition, the system does not rely on precise camera calibration or strict synchronization. The fusion framework is designed to tolerate both geometric misalignment and temporal offset. As shown in Table IV, ablation studies under various calibration conditions and randomized non-reference input orders confirm that the overall reconstruction quality remains stable. Small changes in t-SSIM are mainly due to image cropping rather than alignment failure.
TABLE IV: t-SSIM comparison under different dual-camera deployment settings
t-SSIM MEF setup with DCS
22’ECCV 22’MM 23’ICCV 24’ECCV
HDRTrans DomainPlus SCTNet SAFNet Ours
ori 0.8923 0.8991 0.8868 0.8851 0.9071
w/o calibration 0.8922 0.8994 0.8872 0.8861 0.9079
random non-reference 0.8926 0.8997 0.8842 0.8848 0.9072
lower frame rate 0.8924 0.8985 0.8871 0.8836 0.9066
TABLE V: The ablation study of pre-alignment subnetwork and asymmetric cross-feature fusion subnetwork on Prabhakar’s datasets
pre align fusion PSNR-$\mu$ PSNR-L SSIM-$\mu$ SSIM-L
GLA EFSM ACA $f^{guide}$
$\surd$ $\surd$ 41.29 39.47 0.9728 0.9888
$\surd$ $\surd$ $\surd$ 41.43 39.73 0.9722 0.9890
$\surd$ $\surd$ $\surd$ 41.17 39.33 0.9726 0.9780
$\surd$ $\surd$ $\surd$ $\surd$ 41.80 40.14 0.9731 0.9895
$\surd$ $\surd$ CA $\surd$ 41.61 39.80 0.9726 0.9892
$\surd$ $\surd$ $\surd$ 41.48 39.79 0.9723 0.9893
$\surd$ $\surd$ 41.13 39.12 0.9708 0.9887
40.81 38.70 0.9703 0.9784
Figure 12: Visual comparison of ablation on self-captured videos. The integration of GLA and EFSM influences the overall luminance and jointly reduces ghosting while enhancing detail in dark regions. The luminance histograms show that our method increases the visibility of extremely dark areas without causing an increase in high-luminance pixels.
V-C Ablation study
In this section, we investigate different components of our system individually to understand the capability of the system and the source of efficiency. Specifically, all quantitative evaluations are performed on Prabhakar’s dataset[30].
Global Luminance Alignment. We alter the composition of the input channels and configure four scenarios to investigate the impact of GLA. As shown in Table VI, the first scenario (i.e., our model) exhibits the best performance. The visualization results of the ablation experiments on GLA with Prabhakar’s dataset are shown in Fig. 13, in the first row, other variants fail to restore the overexposed details of the window frame. In the second row, ghosting occurs in the overexposed areas outside the window. The introduction of the global luminance alignment method leads to enhanced detail preservation while effectively mitigating undesirable artifacts.
Figure 13: The contributions of global luminance alignment. Our method effectively suppresses ghosting caused by misalignment and restores more details.
TABLE VI: Ablation investigation about global luminance alignment.
Channels Original
gamma
correction
GLA PSNR-$\mu$ PSNR-$l$ SSIM-$\mu$ SSIM-$l$
9 $\surd$ $\surd$ $\surd$ 41.80 40.14 0.9731 0.9895
6 $\surd$ $\surd$ 41.17 39.33 0.9726 0.9780
6 $\surd$ $\surd$ 41.08 38.97 0.9729 0.9884
6 $\surd$ $\surd$ 41.19 39.54 0.9729 0.9882
Exposure-Guided Feature Selection Module. Table V presents the effectiveness of the EFSM. As shown in the first four rows, without global luminance alignment, EFSM not only fails to achieve the expected improvement but also makes a negative optimization in some cases. This can be attributed to the lack of consistent brightness information during the feature selection process, which may lead to a loss of image details. After introducing GLA, EFSM effectively harmonizes the brightness information, mitigating errors caused by brightness variations. As shown in Fig. 12, this integration results in a significant enhancement in detail preservation while the PSNR-$\mu$ is increased by 0.37 dB over the baseline. The luminance histograms show that the joint role of GLA and EFSM increases the visibility of extremely dark areas without causing an increase in high-luminance pixels. However, in the absence of ESFM, while some improvements are observed in intra-dataset performance (Table V), introducing GLA alone may result in overfitting and suppressing details in high-exposure images. This results in loss of details in dark areas and an overall decrease in image luminance.
We find that feature selection is notably affected by two modulation parameters: feature-guide modulation coefficients and exposure-guide modulation coefficients. To explore this effect more comprehensively, we conduct an in-depth analysis of how varying the modulation coefficient ratios influences performance. As shown in Table VII, we construct five different networks with varying modulation coefficient ratios. Among these, the 2:1 ratio achieves the highest PSNR-$\mu$, indicating optimal performance under this configuration. As shown in Fig. 14, the proposed GLA and EFSM work collaboratively to fully explore exposure information and enhance meaningful details. Specifically, the shadow details in low-exposure features are enhanced, while the overexposed regions in high-exposure features are suppressed.
Figure 14: Visualization of feature maps before and after ESFM. Global luminance alignment is crucial for ESFM, suppressing overexposed areas in the high-exposure input and enhancing underexposed regions in the low-exposure input.
TABLE VII: Experiments on changing the proportions in exposure-guided feature selection module
Proportions PSNR-$\mu$ PSNR-$l$ SSIM-$\mu$ SSIM-$l$
1:2 41.41 39.66 0.97281 0.97907
1:1 41.43 39.70 0.97284 0.98918
2:1 41.80 40.14 0.97311 0.98953
4:1 41.68 40.00 0.97251 0.98934
8:1 41.72 40.02 0.97255 0.98938
Asymmetric Cross-feature Fusion Subnetwork. As shown in Table V, replacing the asymmetric cross-feature fusion subnetwork with a generally used attention subnetwork [36] (line 7) results in a 0.67 dB drop in PSNR-$\mu$. Fig. 16 and Fig. 15 further confirms that our method effectively mitigates misalignment artifacts. Additionally, we evaluate the role of the asymmetric cross-attention (ACA) by comparing it with cross-attention (CA) mechanism. The ACA improves PSNR-$\mu$ by 0.19 dB (line 5). The cross-scale guidance features $f^{guide}$ further enhance fusion accuracy, reducing artifacts and contributing to a 0.32 dB gain in PSNR-$\mu$.
Figure 15: Visual comparison of asymmetric cross-attention and cross-attention. Our method reduces misalignment artifacts.
Restoration Subnetwork In this section, we conduct ablation experiments to evaluate the role of the frequency-domain design and multiscale structure in HDR reconstruction. As shown in Table VIII, the baseline model (merging network in AHDNet) achieves a PSNR-$\mu$ of 40.81 and an SSIM-$\mu$ of 0.9719. First, we remove the frequency-domain design, and the results show that the PSNR-$\mu$ dropped by 0.09. Next, we perform an ablation experiment on the multiscale structure by removing it. The results indicate that, without the multiscale structure, the PSNR-$\mu$ decreased by 0.22. Our full model, which incorporates both the frequency-domain design and multiscale structure, yields the highest performance, achieving a PSNR-$\mu$ of 41.80 and an SSIM-$\mu$ of 0.97311.
TABLE VIII: Visual comparison of the separate effects of restoration network.
model PSNR-µ PSNR-L SSIM-µ SSIM-L
baseline 40.81 38.60 0.97192 0.98813
w/o mutil-scale 41.58 39.86 0.97249 0.98911
w/o DWT 41.71 40.04 0.97286 0.98937
Ours 41.80 40.14 0.97311 0.98953
Figure 16: Visual comparison of ablation on self-captured videos. Our method reduces most ghosting artifacts.
VI CONCLUSION
In this work, we propose a dual-camera system (DCS) to mitigate the temporal inconsistency with AE methods, and existing MEF methods can be seamlessly extended to our system with temporally consistent output. In addition, we introduce an exposure adaptive fusion network with two novel subnetworks to provide a more stable and efficient solution for video HDR reconstruction. Experimental results demonstrate that our approach outperforms existing methods, offering stable performance and high-quality reconstruction in complex real-world video environments. Our dual-camera framework provides a promising direction for real-time HDR video capture.

### v2

@@
 Abstract
-In HDR video reconstruction, exposure fluctuations in reference images from alternating exposure methods often result in flickering.
-To address this issue, we propose a dual-camera system (DCS) for HDR video acquisition, where one camera is assigned to capture consistent reference sequences, while the other is assigned to capture non-reference sequences for information supplementation.
-To tackle the challenges posed by video data, we introduce an exposure-adaptive fusion network (EAFNet) to achieve more robust results.
-EAFNet introduced a pre-alignment subnetwork to explore the influence of exposure, selectively emphasizing the valuable features across different exposure levels.
-Then, the enhanced features are fused by the asymmetric cross-feature fusion subnetwork, which explores reference-dominated attention maps to improve image fusion by aligning cross-scale features and performing cross-feature fusion.
-Finally, the reconstruction subnetwork adopts a DWT-based multiscale architecture to reduce ghosting artifacts and refine features at different resolutions.
-Extensive experimental evaluations demonstrate that the proposed method achieves state-of-the-art performance on different datasets, validating the great potential of the DCS in HDR video reconstruction.
-The codes and data captured by DCS will be available at https://github.com/zqqqyu/DCS.
+High Dynamic Range (HDR) video acquisition using the alternating exposure (AE) paradigm has garnered significant attention due to its cost-effectiveness with a single consumer camera.
+However, despite progress driven by deep neural networks, these methods remain prone to temporal flicker in real-world applications due to inter-frame exposure inconsistencies.
+To address this challenge while maintaining the cost-effectiveness of the AE paradigm, we propose a novel learning-based HDR video generation solution.
+Specifically, we propose a dual-stream HDR video generation paradigm that decouples temporal luminance anchoring from exposure-variant detail reconstruction, overcoming the inherent limitations of the AE paradigm.
+To support this, we design an asynchronous dual-camera system (DCS), which enables independent exposure control across two cameras, eliminating the need for synchronization typically required in traditional multi-camera setups.
+Furthermore, an exposure-adaptive fusion network (EAFNet) is formulated for the DCS system.
+EAFNet integrates a pre-alignment subnetwork that aligns features across varying exposures, ensuring robust feature extraction for subsequent fusion, an asymmetric cross-feature fusion subnetwork that emphasizes reference-based attention to effectively merge these features across exposures, and a reconstruction subnetwork to mitigate ghosting artifacts and preserve fine details.
+Extensive experimental evaluations demonstrate that the proposed method achieves state-of-the-art performance across various datasets, showing the remarkable potential of our solution in HDR video reconstruction.
+The codes and data captured by DCS will be available at https://zqqqyu.github.io/DCS-HDR/.
 Index Terms:
-HDR video, dual-camera system, multi-exposure image fusion
+HDR video, multi-exposure image fusion, luminance alignment
 I Introduction
-Capturing the full range of illumination in high dynamic range (HDR) scenarios is a challenging task for standard digital cameras [1, 2].
-Thanks to the multi-exposure fusion (MEF) technology [3, 4, 5], we can successfully capture HDR images in static scenarios by fusing multiple images of different exposure times.
-However, capturing HDR videos with a standard digital camera in dynamic scenes has not yet been effectively resolved.
-Early studies concentrated on designing hardware systems for directly capturing HDR video, including internal/external beam splitters [6, 7, 8], specialized sensors [9, 10], modulo cameras [11] and neuromorphic cameras [12], etc.
+Capturing the full range of illumination in high dynamic range (HDR) scenarios is a challenging task for standard digital cameras [10, 18, 29].
+Thanks to the multi-exposure fusion (MEF) technology [1, 45], we can successfully capture HDR images in static scenarios by fusing multiple images of different exposure times.
+However, capturing HDR videos with a standard digital camera in dynamic scenes has not yet been effectively resolved [51, 50].
+Early studies concentrated on designing hardware systems for directly capturing HDR video, including internal/external beam splitters [28, 39, 47], specialized sensors [7, 15, 20, 38], modular cameras [61] and neuromorphic cameras [14], etc.
 Although these solutions show good potential for capturing high-quality HDR videos, their high cost and bulky design severely limit their widespread adoption.
-Figure 1: Qualitative comparison between SAFNet and AE methods.
-Existing AE methods result in flickering artifacts (luminance fluctuations exceeding 30).
-In contrast, our system ensures consistent illumination across reference frames, enabling temporally consistent HDR reconstruction results (luminance fluctuations limited to under 1).
-To address the limitations of these issues, researchers [13, 14, 15] have turned to computational methods that leverage inexpensive standard digital cameras, providing a promising direction for HDR video reconstruction.
-With the rapid advancement of deep learning, researchers make efforts to solve HDR imaging problems in dynamic scenes with deep neural networks.
-These methods can be broadly categorized into two types.
-One approach follows the MEF attempts, aiming to reconstruct an HDR image by fusing misaligned multi-exposure images.
-They typically use a specific exposure image as a reference frame for HDR reconstruction.
-As a result, capturing HDR video using these methods often requires a large number of additional exposure images, making it difficult to meet the real-time requirements of video capture.
-To overcome this limitation, alternating exposure (AE) methods [16, 17, 18, 19, 20] have been proposed, which allow varying exposures as reference frames and eliminate the need for additional exposure images.
-However, aligning the brightness of reference frames of different exposures and achieving consistent and stable lighting in the reconstruction video is still challenging.
-Although existing AE methods have achieved excellent performance on public benchmarks, they struggle to handle complex real-world scenes and lighting variations (shown in Fig.
-1).
-Generally, both existing MEF methods and AE methods struggle to achieve high-quality video capture in dynamic scenes.
-Compared to AE methods, MEF methods could produce HDR image sequences with stable and consistent lighting because of their use of a fixed-exposure reference.
-Noticing the wide use of multi-camera system in consumer electronic devices, it is possible to introduce multi-camera system for high-quality HDR video capture.
-Based on these insights, we develop a dual-camera system (DCS) for capturing high-quality HDR videos.
-In the DCS, one camera is assigned to capture reference images, while the other is assigned to capture non-reference images for MEF.
-In this way, the DCS could effectively mitigate the temporal inconsistencies associated with AE methods while avoiding the acquisition latency inherent in MEF.
-Within DCS, HDR video reconstruction with MEF faces greater challenges than static images for two main reasons.
-First, exposures of input frames become uncertain.
-Second, the backgrounds from non-reference images are no longer aligned with the reference frame.
-Existing MEF methods can be categorized into two main approaches: optical-flow-based and attention-based.
-The former relies on optical flow to align inputs before fusion, which often suffers from a low accuracy in estimating optical flow.
-On the other hand, attention-based methods, while offering greater flexibility, mainly depend on the quality of the attention maps for fusion.
-Therefore, existing attention-based MEF methods encounter two main issues in reconstructing HDR videos within DCS.
-First, they fail to adequately leverage exposure information, relying primarily on texture and luminance to calculate attention maps, which may suffer from the various exposure settings.
-Second, they hardly consider the importance of reference frames, particularly in the computation of attention maps, where equal weights are often assigned to both reference and non-reference frames.
-This leads to attention being skewed by non-reference inputs, especially in complex, dynamic scenes, thereby hindering the effective elimination of artifacts.
-Against the above limitations, we propose an exposure-adaptive fusion network (EAFNet) to achieve robust and generalized HDR reconstruction.
+Figure 1: Overview of challenges in the alternating exposure (AE) paradigm and our proposed dual-stream paradigm for stable HDR video generation.
+(a) existing AE paradigm fuses alternating exposures (LE/ME/HE) with varying reference frames, resulting in flickering artifacts (luminance fluctuations exceeding 30).
+(b-d) Our dual-stream paradigm employs a fixed-exposure reference sequence for stable luminance anchoring, while a separate stream handles exposure-variant enhancement, ensuring consistent illumination across reference frames, enabling temporally consistent HDR reconstruction results (luminance fluctuations limited to under 1).
+The alternating exposure (AE) paradigm, first introduced by Kang et al.
+[25], enables low-cost HDR video acquisition using a single camera with rapidly switching exposures, but it also introduces challenges in maintaining temporal consistency and suppressing ghosting artifacts.
+Traditional AE-based methods adopt strategies such as motion-compensated fusion [24, 13] and optical flow-based reconstruction [35] to align and merge complementary exposures for HDR reconstruction, but they remain highly sensitive to motion misalignment, struggle with occlusions and non-rigid deformations, and often rely on hand-crafted heuristics that lack generalization.
+Recent deep learning approaches [21, 9, 54] have advanced this direction by employing deep neural architectures with stronger feature representations and learning-based alignment mechanisms.
+These methods have achieved promising results by training on synthetic datasets with well-structured exposure patterns and often produce temporally consistent outputs under controlled settings.
+However, in real-world scenarios, variations in brightness distribution caused by changes in scene content, illumination, or sensor response can disrupt the learned alignment and fusion behavior.
+As a result, networks that perform well in ideal conditions often suffer from temporal flickering or inconsistent reconstruction when deployed in real environments (see Fig.
+1(d)).
+This instability highlights a fundamental limitation of the AE paradigm itself, rather than just model design, thus leading us to explore a new solution that circumvents inter-frame exposure inconsistency.
+Specifically, in the AE paradigm, detail selection and global luminance determination are inherently entangled at the input level.
+Existing methods [9, 54] typically process both aspects within a single modeling stream, without explicit disentanglement in input design or feature representation.
+As a result, the network must simultaneously decide which regions to source from each exposure and how to establish the overall luminance mapping within the same representational space.
+As shown in Fig.
+1, this tight coupling leads to feature interference and makes the learned mapping strongly dependent on the exposure patterns seen during training, thereby reducing robustness and temporal stability under distribution shifts.
+To address this, we propose a dual-stream generation paradigm that decouples temporal luminance anchoring from exposure-variant detail reconstruction, thereby reducing feature interference and enhancing generalization across varying exposure patterns and scene conditions.
+Our paradigm is meticulously designed at both the hardware system and algorithmic levels, forming a complete HDR video generation solution that supports practical deployment and robust reconstruction in real-world environments.
+Our hardware architecture is designed to implement an asynchronous dual-camera system (DCS) that capitalizes on the prevalent availability of multi-camera hardware in consumer electronic devices.
+Unlike prior dual-camera HDR imaging systems [11, 31] designed for static scenes under strict synchronization constraints, our DCS is tailored for dynamic video scenarios.
+By relaxing the synchronization requirement, it supports independent exposure control across two cameras.
+This design not only facilitates flexible and practical deployment without precision alignment but also avoids the frame rate bottleneck imposed by long-exposure frames in synchronized systems.
+As a result, our system enables high-quality HDR video capture at high frame rates.
+Specifically, the designed DCS system utilizes a fixed-exposure reference sequence as the temporal baseline, while simultaneously leveraging a varying-exposure stream to enrich the dynamic range.
+This system-level approach addresses key limitations of existing video methods.
+The reference stream mitigates the temporal inconsistencies and flickering commonly observed in AE-based single-camera systems, while the division of labor between the two cameras eliminates the need for precise synchronization inherent in traditional multi-camera systems, allowing for independent and flexible exposure control.
+Furthermore, the proposed DCS remains fully compatible with most existing image HDR deghosting algorithms, enabling temporally consistent and robust HDR video reconstruction.
+Building on the proposed paradigm and DCS system, we now turn to the algorithmic level.
+We first revisit existing HDR deghosting approaches to extract insights and expose their limitations.
+Current methods can be broadly divided into optical-flow-based [44, 42] and attention-based [56, 34] strategies.
+The former relies on optical flow to align inputs before fusion, which often suffers from low accuracy in estimating optical flow.
+On the other hand, attention-based methods, while offering greater flexibility, their attention maps are typically computed only from texture and luminance cues, without explicitly leveraging exposure information.
+Moreover, they generally assign equal importance to reference and non-reference frames, which allows attention to be skewed by non-reference inputs in complex scenes.
+In our DCS, the presence of parallax introduces additional variations between inputs, making it even more important to design attention mechanisms that emphasize reference frames.
+Consequently, both categories exhibit limited robustness in dynamic, exposure-varying, and parallax-affected settings, where inaccurate alignment and exposure-insensitive attention lead to ghosting and color distortion.
+To address this, we propose an exposure-adaptive fusion network (EAFNet) tailored for the proposed DCS system to achieve robust and generalized HDR reconstruction.
 Our EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a reconstruction subnetwork.
@@
 The asymmetric cross-feature fusion subnetwork aims to explore reference-dominated attention maps, align cross-scale features, and integrate cross-feature information to improve the fusion.
-Additionally, a multiscale reconstruction subnetwork is introduced to suppress ghosting artifacts and enhance fine details.
-Summary, our contributions can be summarized as follows:
-  • We propose a new HDR video generation paradigm that uses a fixed medium-exposure sequence as the temporal reference, while leveraging a second varying-exposure stream to enrich dynamic range.
+Additionally, a reconstruction subnetwork is introduced to suppress ghosting artifacts and enhance fine details.
+In summary, our contributions can be summarized as follows:
+  • Paradigm: We introduce a dual-stream HDR video generation paradigm that explicitly decouples temporal luminance anchoring from exposure-variant detail reconstruction.
+Our approach employs a fixed-exposure stream to maintain temporal alignment across frames, while a complementary stream with varying exposures enhances the dynamic range.
 This design fundamentally improves temporal consistency and reconstruction stability.
-  • We designed and implemented an asynchronous dual-camera system to validate the feasibility of our proposed paradigm and bridge the gap between algorithmic design and practical deployment.
-Furthermore, existing MEF methods can be seamlessly applied within our system, achieving temporally consistent reconstruction.
-  • Built upon MEF methods, we propose a novel model design, EAFNet.
-The proposed pre-alignment subnetwork leverages exposure information and explores the intrinsic properties of the images, ensuring that the most relevant features receive attention despite varying exposure levels.
-An asymmetric cross-feature fusion subnetwork is introduced to explore reference-dominated attention maps, achieving smoother and more consistent fusion through cross-scale alignment and cross-feature interaction.
-  • Our EAFNet achieves state-of-the-art results on public benchmarks. Both qualitative and quantitative experiments demonstrate the effectiveness of our EAFNet and DCS.
+  • System: We design and implement an asynchronous dual-camera system to validate the feasibility of our proposed solution and bridge the gap between algorithmic design and practical deployment.
+Unlike traditional synchronized setups constrained by long-exposure frames, our system enables high-frame-rate video capture in dynamic scenes by supporting independent exposure control without requiring hardware-level synchronization.
+Moreover, the system seamlessly integrates with existing image deghosting methods to achieve temporally consistent video reconstruction.
+  • Method: To support our dual-camera system, we propose a novel model design, EAFNet.
+We leverage exposure information and explore the intrinsic properties of the images, ensuring that the most relevant features receive attention under different exposure conditions.
+Furthermore, we explore reference-dominated attention and effectively fuse complementary features across exposures, enabling robust representation under diverse lighting scenarios.
+  • Results: Our proposed dual-stream HDR video generation paradigm demonstrates significant advancements across multiple aspects.
+Through extensive qualitative and quantitative experiments, we validate the effectiveness of our contribution.
+The rest of this article is organized as follows.
+Section II presents the related works.
+Section III introduces our solution for video HDR capturing and our dual-camera system.
+Section IV introduces the proposed EAFNet in detail.
+Section V presents the experimental results, and the conclusion is given in Section VI.
 II RELATED WORK
 II-A HDR Image Reconstruction.
-II-A1 Traditional Methods.
-Traditional multi-exposure image fusion HDR imaging methods usually solve the ghosting problem in the image domain, and currently there are two main types of solutions:
-Motion-aligned-based.
-These methods focus on a global alignment of the LDR input and then reject the unaligned pixels before image fusion to achieve de-ghosting.
-Bogoni et al.
-[4] used optical flow to estimate motion vectors between LDR images and warp the moving pixels accordingly.
-Kang et al.
-[16] improved the alignment results by using the exposure time information to transform the intensity of LDR images to the brightness domain before aligning them with the optical flow method.
-Sen et al.
-[5] proposed a patch-based energy minimization method that integrates alignment and HDR reconstruction into a joint optimization.
-Hu et al.
-[21] optimized alignment using luminance and gradient consistency in the transform domain.
-Hafer et al.
-[22] proposed an energy minimization approach to jointly compute HDR radiance and displacement fields.
-These motion-alignment methods heavily rely on the accuracy of the alignment algorithms.
-Motion-rejection-based.
-Motion-rejection methods detect and exclude motion regions before fusion.
+II-A1 Traditional Methods
+Traditional multi-exposure HDR image fusion methods address the ghosting problem using two main approaches: motion-aligned-based and motion-rejection-based methods.
+Motion-aligned methods focus on globally aligning LDR images, followed by rejection of unaligned pixels to eliminate ghosting.
+Early works by Bogoni et al.
+[1] used optical flow for motion estimation, while later improvements by Kang et al.
+[25] incorporated exposure time information for enhanced alignment.
+Further advancements include patch-based energy minimization by Sen et al.
+[45] and luminance-based optimization by Hu et al.
+[17].
+Motion-rejection methods, on the other hand, identify and exclude motion regions before fusion.
 Grosch et al.
-[23] identified motion regions based on intensity differences, while Jacobs et al.
-[13] used weighted variance to detect ghosting.
-He et al.
-[24] refined motion detection by combining joint probability density with graph-cut-based energy minimization.
-Zhang et al.
-[25] removed motion pixels using image gradient information, while Lee et al.
-[26] and Oh et al.
-[27] introduced rank minimization to detect motion and reconstruct HDR images.
-However, discarding motion regions often compromises information recovery in static areas.
-II-A2 Deep Learning Approaches.
-With the rapid advancement of deep learning, neural network-based methods for HDR image reconstruction have gained significant attention and have emerged as the prevailing approach in the field, these can be categorized primarily into the following types:
+[12] and Jacobs et al.
+[19] detected motion using intensity differences and weighted variance, respectively, while He et al.
+[16] and Zhang et al.
+[59] refined motion detection with graph-cut and image gradient techniques.
+II-A2 Deep Learning Approaches
+With the rapid advancement of deep learning, neural network-based methods for HDR image reconstruction have gained significant attention and have emerged as the prevailing approach in the field.
+These can be categorized primarily into the following types:
 Flow-based Alignment.
@@
 [53] introduced a flow-free approach, treating HDR imaging as an image transformation problem.
-They first applied homography to align multiple frames globally and then used a multiscale network to map LDR images to HDR images directly.
 A similar approach was also employed in [57].
@@
 Attention-based Alignment.
-AHDRNet [36] introduced the attention mechanism into multi-exposure HDR reconstruction for dynamic scenes, coupling it with a densely connected dilated residual module to achieve alignment and fusion.
+AHDRNet [56] first introduced the attention mechanism into multi-exposure HDR reconstruction for dynamic scenes, coupling it with a densely connected dilated residual module to achieve alignment and fusion.
+Subsequent works focused on refining alignment strategies.
 Liu et al.
-[37] introduced a Pyramid Cascading Deformable (PCD) module, using multiscale deformable convolutions for feature alignment in LDR images.
+[33] introduced a Pyramid Cascading Deformable (PCD) module, using multi-scale deformable convolutions for feature alignment in LDR images.
 Chen et al.
-[38] proposed novel attention mechanisms and progressive texture fusion to improve reconstruction.
-DomainPlus [39] utilized multiscale attention with fronted fusion to perform coarse-to-fine fusion.
-HDR-Trans et al.
-[40] added Transformer to the CNN-based HDR architecture, combining the advantages of CNN in extracting local information and Transformer in capturing global information.
-Chen et al.
-[41] proposed a transformer model for HDR imaging with three steps: alignment, fusion, and reconstruction.
-Yan et al.
-[42] proposed a HyHDRNet consisting of a content alignment subnetwork and a transformer-based fusion subnetwork to improve performance.
-SCTNet [43] integrated spatial and channel attention in a transformer-based network to improve semantic coherence.
-Despite advances in these approaches, they still generate undesirable ghosts in high-contrast regions.
-II-A3 Dual-lens-based system
+[3] and DomainPlus [62] further improved fusion through progressive attention and coarse-to-fine multi-scale designs.
+Transformer-based methods such as HDR-Trans [34], Chen et al.
+[4], and HyHDRNet [55] integrated global context modeling into HDR pipelines, enhancing feature representation.
+SCTNet [46] combined spatial and channel attention to improve semantic consistency.
+Despite advances in these approaches, most existing methods lack explicit modeling of exposure variations, resulting in sensitivity to brightness variations.
+Moreover, treating reference and non-reference frames equally during fusion often leads to ghosting artifacts, particularly in high-contrast regions dominated by non-reference content.
+Dual-lens-based system.
 Dual-lens HDR imaging leverages paired images with different exposures to synthesize high-quality HDR content.
 Dong et al.
-[44] employ a soft-warping step to align long-exposure features to the short-exposure reference, Li et al.
-[45] propose a dual-stage model: a guided 3D exposure correction network estimates a coarse HDR image from the short exposure input, followed by a guided denoising transformer that refines the result using long-exposure features.
+[11] and Li et al.
+[31] align features under the assumption of simple 1D parallax between rectified stereo pairs.
 While effective under controlled conditions, both frameworks assume the only misalignment present is 1D parallax between image pairs.
 This assumption is embedded in their data collection and rectification procedures, simplifying alignment to left–right disparities but limiting their ability to model more complex distortions.
-In dynamic scenes, exposure time mismatch introduces additional misalignments: high-exposure frames exhibit motion blur not present in low-exposure frames.
+In dynamic scenes, mismatched exposure times introduce complex misalignments, including motion blur in long-exposure frames that cannot be captured by disparity-based alignment alone.
 II-B HDR Video Reconstruction
@@
 Traditional Methods.
-Similar to multi-exposure HDR image fusion, most methods utilize alternating exposure sequences to align multiple frames for HDR video reconstruction.
+Most HDR video reconstruction methods utilize alternating exposure sequences to align multiple frames for HDR video reconstruction.
 Kang et al.
 [25] introduced the first algorithm in this category, employing global and local registration to align multi-exposure images at each time step and merge them into a single HDR frame.
-Kalantari et al.
-[17] employed patch-based optimization to reconstruct missing images from different exposures.
 Mangiat et al.
@@
 LAN-HDR [8] aligned adjacent frames with the reference frame using luminance-based attention scores.
-HDRFlow [51] designed an efficient flow network with multiscale large kernels to handle large motions.
+HDRFlow [54] designed an efficient flow network with multi-scale large kernels to handle large motions.
 Cui et al.
 [9] introduced exposure completion for neural HDR rendering, using feature interpolation and a novel decoding architecture to improve HDR video reconstruction.
-Nonetheless, the varying exposures of the reference frames create a significant challenge in maintaining temporal consistency in video reconstruction.
+Nonetheless, alternating exposure setups inherently introduce brightness inconsistencies across reference frames, leading to temporal instability in the reconstructed video.
+In contrast, our design adopts a fixed-exposure reference strategy, which naturally preserves brightness consistency and mitigates flickering artifacts.
 Figure 2: Visualization of our dual-camera system.
 The primary camera captures continuous medium-exposure sequences as reference for temporal consistency, while the secondary camera alternates between low- and high-exposure to provide complementary information for reconstruction.
-III DUAL-CAMERA SYSTEM FOR HDR VIDEO
+III DUAL-CAMERA SYSTEM FOR CAPTURING HDR VIDEO
 Overview.
-A mainstream solution for HDR video reconstruction is generating video from image sequences captured at alternating exposure levels.
+An inexpensive mainstream solution for HDR video reconstruction is generating video from image sequences captured at alternating exposure levels [21, 9, 54].
 However, this approach inherits the exposure limitations of the reference frame, often resulting in visible artifacts and temporal flickering in the generated video, significantly constraining its practical application.
-To address this, we propose a dual-camera system (As shown in Fig.
-2 (b)), where the primary camera continuously captures medium-exposure images to ensure temporal consistency, while the secondary camera captures low-exposure and high-exposure images to provide complementary information.
+To address this, we propose a dual-stream paradigm supported by a dual-camera system (as shown in Fig.
+2), where the primary camera continuously captures medium-exposure frames to ensure temporal consistency, while the secondary camera captures low-exposure and high-exposure frames to provide complementary information.
 III-A Dual-Camera System
@@
 This reduces potential errors and improves the subsequent processing steps.
-Video Capture.
-Both cameras are triggered to start capturing images simultaneously, but subsequent image capture can occur asynchronously.
-One camera captures the medium-exposure reference sequence $L_{reference}(t)$, while the other captures the alternating exposure non-reference sequence $L_{non-reference}(t)$ with low and high exposure.
-Both cameras are set to the same resolution of $w\times h$.
-Due to different exposure time settings, the videos captured by the two cameras are not timestamp-synchronized.
-However, our method does not require precise synchronization.
-We use timestamp information to roughly group the data, where each group consists of frames from $L_{reference}(t)$ and two high- and low-exposure images from $L_{non-reference}(t)$ (e.g., with exposure times of (-2, 0, 2 EV), every four frames from the reference sequence, along with one pair of high- and low-exposure images, form a single group).
-Each frame, along with the two high- and low-exposure images, is used as input to the network.
-The final reconstructed HDR video matches the frame rate of $L_{reference}(t)$.
-To ensure a fair comparison, we capture test datasets for both AE and MEF using our system.
-Detailed information can be found in Sec.
-V-B, and the datasets will be available.
-Essentially, our system can be viewed as a multi-exposure HDR fusion task for dynamic scenes, where luminance alignment between neighboring frames is unnecessary due to consistently exposed reference frames.
-Existing MEF methods can be seamlessly extended to achieve consistent exposure results.
-To address challenges from video data and the variability of real-world scenes, we propose SAFNet, a more robust network, detailed in Section IV.
+Video Capture and System Design.
+Our dual-camera system employs two identical cameras with resolution $w\times h$.
+In practical operation, the cameras are run asynchronously: one continuously records a medium-exposure reference sequence $L_{ref}(t_{1})$ to provide temporally consistent baseline frames, while the other alternates between low and high exposures $L_{non\mbox{-}ref}(t_{2})$ to supply the luminance diversity required for dynamic range expansion.
+To leverage the information from two sequences, we construct approximate pairings using timestamp metadata: as shown in Fig.
+2(a), for each low/high exposure pair in the alternating stream, we gather reference frames falling within its defined temporal neighborhood and combine each of them with the same low/high pair to form input groups, enabling the reuse of one low/high pair across multiple reference frames.
+This pairing does not require frame-level precision, as the scenes are inherently dynamic and minor temporal offsets have a negligible impact on multi-exposure fusion.
+Each input group is then fed into the network to reconstruct HDR video at the same frame rate as $L_{ref}(t_{1})$.
+Overall, the design formulates the task as multi-exposure HDR fusion for dynamic scenes, leveraging the consistent exposure of the reference sequence to eliminate the need for luminance alignment between neighboring frames.
+Existing HDR image deghosting methods can be seamlessly extended to achieve consistent luminance results.
+To address challenges from video data and the variability of real-world scenes, we propose EAFNet, a more robust network, detailed in Section IV.
+III-B Further Discussion
+To further clarify our contributions, we discuss the proposed paradigm from three perspectives:
+III-B1 Relationship with Previous HDR video pipelines
+Traditional HDR video generation methods often rely on specialized sensors or expensive camera setups, such as Quad-Bayer sensors [20], Beam Splitter systems [28], etc.
+These methods produce high-quality HDR output through specialized hardware designs for capturing frames with different exposures, but come with significant hardware costs and are typically limited to high-end applications.
+In contrast, our dual-stream paradigm not only lowers hardware costs but also enhances flexibility, making it feasible for HDR video generation on consumer-grade devices.
+Additionally, many existing methods address hardware limitations by employing alternating exposure techniques [54, 2, 8], where a single consumer-grade camera alternates between different exposure levels to capture HDR video.
+Our paradigm retains the cost-effectiveness of traditional AE methods while also addressing key challenges such as temporal consistency and robustness in real-world dynamic imaging scenarios.
 Figure 3: The architecture of EAFNet consists of a pre-alignment subnetwork, an asymmetric cross-feature fusion subnetwork, and a restoration subnetwork.
@@
 The asymmetric cross-feature fusion subnetwork improves image fusion by aligning cross-scale features and performing cross-feature fusion.
-The reconstruction subnetwork adopts a DWT-based multiscale architecture to reduce ghosting and refine features at different resolutions.
-III-B Discussion.
-The proposed DCS offers several advantages over traditional single-camera AE methods for HDR video reconstruction:
-Stable Lighting Conditions. The DCS provides more stable exposure across frames, avoiding the fluctuations typical in alternating exposure systems.
-Potential in Low-Light Conditions.
-The DCS performs better in low-light environments, where long exposure times in a single-camera setup can lead to artifacts and delays.
-The dual-camera approach mitigates these issues by distributing exposure tasks between the cameras.
-Offline HDR Video Reconstruction.
-The DCS setup supports offline HDR video reconstruction, allowing continuous high-quality video capture without real-time processing, ensuring high-quality capture without real-time processing delays.
-In contrast, AE methods require real-time reconstruction to avoid flickering, which can introduce undesirable artifacts if not processed immediately.
-Independent Automatic Exposure.
-Each camera in the DCS can independently apply established automatic exposure strategies, eliminating the need for recalibration with each frame and simplifying real-time operation.
-In contrast, AE methods require constant adjustments to exposure settings, complicating real-time capture.
-Seamless Integration with Encoding Strategies.
-The dual-camera framework integrates easily with existing video encoding and decoding strategies.
-The secondary camera can capture frames at a low frame rate (e.g., 1fps or lower), minimizing additional bandwidth requirements, thus ensuring efficient video transmission.
+The reconstruction subnetwork adopts a multi-scale architecture to reduce ghosting and refine features at different resolutions.
+III-B2 Relationship with Previous dual-camera systems
+Dual-camera systems have been used in HDR imaging [11, 31], but they typically require strict synchronization, which is challenging in dynamic video scenes due to exposure differences between cameras, leading to ghosting artifacts.
+In contrast, our system performs exposure fusion without synchronization, enabling better adaptability to dynamic environments.
+Additionally, while synchronized systems are constrained by long exposure times that lower frame rates, our system decouples temporal stability from exposure variation, allowing flexible exposure settings and maintaining high frame rates.
+The reference stream controls the frame rate, making our system particularly suited for high-frame-rate HDR video capture, especially in low-light conditions.
+Beyond HDR imaging, dual-camera frameworks have also been explored in related video enhancement tasks.
+For instance, Cheng et al.
+[5] and H2-Stereo [6] proposed dual-camera super-resolution systems that pair a high-resolution low-frame-rate stream with a low-resolution high-frame-rate stream to reconstruct videos with high spatiotemporal resolution.
+Similarly, dual-camera designs have been applied in video de-moiré [49].
+While these methods also exploit asymmetric information across cameras, their optimization targets are fundamentally different from ours: they generally assume consistent exposure between cameras and focus on improving spatial or temporal fidelity, rather than expanding dynamic range.
+In contrast, our framework introduces an exposure-asymmetric dual-camera configuration: a reference camera maintains a stable mid-level exposure to anchor temporal consistency, while an auxiliary camera alternates between low and high exposures to provide extreme luminance details.
+This design effectively decouples temporal stability from exposure diversity, a direction not explored in prior work.
+Furthermore, we relax the requirement for frame-level synchronization by adopting group-wise asynchronous fusion, where multiple reference frames are paired with an alternating-exposure pair from the auxiliary camera and jointly processed by a dedicated deghosting network.
+This pipeline enables high-frame-rate HDR video capture with minimal flicker, a capability not achievable with single-camera alternating exposure or conventional dual-camera systems designed for other tasks.
+In addition, our system exhibits high scalability and strong generalization capability.
+It does not rely on device-specific calibration or training data.
+Instead, multi-exposure fusion models trained on publicly available datasets can be seamlessly integrated into our framework without any retraining or fine-tuning.
+Leveraging this flexibility, our system can directly produce flicker-free HDR videos from real-world captures.
+In contrast, existing dual-camera HDR solutions typically require carefully calibrated camera pairs and task-specific training, which significantly restricts their deployment in practical scenarios.
+To the best of our knowledge, none of the prior methods provide such plug-and-play adaptability while maintaining high temporal stability.
+III-B3 The advantages of the proposed video generation paradigm
+The proposed video generation paradigm offers several distinct advantages.
+First, the dual-stream design introduces an exposure-asymmetric configuration that explicitly decouples temporal stability from exposure diversity.
+This principle alleviates the feature interference inherent in AE-based HDR systems and improves robustness in dynamic real-world environments.
+Second, the paradigm offers flexibility in exposure scheduling and frame rate control.
+By anchoring the output frame rate to the reference stream, the system can deliver high-frame-rate HDR video while simultaneously leveraging the auxiliary stream for extended dynamic range, which is particularly beneficial in low-light or high-motion conditions.
+In this asynchronous dual-stream capture scheme, the reference stream follows standard auto-exposure to provide a stable baseline, while the auxiliary stream operates at a lower frame rate with fixed or preset exposures.
+Such a design improves energy efficiency and reduces bandwidth and storage overhead, enabling seamless integration with standard video encoding and transmission protocols.
+Third, the paradigm achieves strong scalability and generalization.
+Multi-exposure fusion models trained on publicly available datasets can be directly applied without device-specific calibration or retraining, enabling plug-and-play deployment.
+Finally, the design is highly practical for integration into consumer-grade devices.
+Existing multi-camera systems can readily adopt the dual-stream setup without the need for specialized sensors or costly calibration procedures, broadening the accessibility of HDR video capture.
 IV METHOD
 IV-A Overview
-Given a set of LDR images of a dynamic scene $\left\{L_{1},L_{2},\ldots,L_{n}\right\}$ with different exposures, HDR imaging aims to generate an HDR image $H$ aligned to a prescribed reference image.
-Fig.
-3 shows an overview of the proposed EAFNet.
-Our proposed EAFNet consists of three main subnetworks: the pre-alignment subnetwork, the asymmetric cross-feature fusion subnetwork, and the restoration subnetwork.
-IV-B Pre-alignment
+In this section, we introduce the overall framework of the proposed EAFNet, built upon a pre-alignment subnetwork (in Sec.
+IV-B), an asymmetric cross-feature fusion subnetwork (in Sec.
+IV-C) and a restoration subnetwork (in Sec.
+IV-D).
+As shown in Fig.
+3, the pre-alignment subnetwork aims to improve detail preservation and color accuracy by exploring exposure information and extracting meaningful features from the non-reference image.
+It incorporates Global Luminance Alignment (GLA) to ensure consistent brightness across varying exposures, and an Exposure-guide Feature Selection Module (EFSM) to identify and focus on the most relevant regions based on the exposure information.
+These mechanisms work together to enhance the input features before fusion.
+Next, the asymmetric cross-feature fusion subnetwork plays a critical role in blending features from different exposures while emphasizing the most important details.
+It processes the low- and high-exposure inputs through separate asymmetric cross-feature fusion blocks, each built with a multi-scale design.
+The fusion is driven by an Asymmetric Cross-Attention (ACA), which leverages reference-dominated attention maps to improve the alignment of query and key features, facilitating more effective feature propagation across multiple scales.
+Finally, the restoration subnetwork addresses common issues such as ghosting artifacts.
+It is designed to suppress ghosting while preserving fine details at multiple scales.
+IV-B pre-alignment subnetwork
 The pre-alignment subnetwork aims to extract meaningful details from the non-reference image by leveraging exposure information.
-Global luminance alignment (GLA) is introduced to maintain consistent luminance across varying exposures, while the exposure-guide feature selection module (EFSM) explores the correlation between inputs and exposure to focus on the most relevant regions for fusion.
-Global Luminance Alignment.
-Following [53, 36], we employ three LDR images $\left\{L_{l},L_{m},L_{h}\right\}$ (sorted by their exposure length) as input.
-The images first undergo gamma correction [53], defined as:
-$H_{i}=L_{i}^{\gamma}/{t_{i}},\ \ i=l,m,h$ (1)
-where $\gamma>1$ denotes the gamma correction parameter and $t_{i}$ denotes the exposure time.
-The indices $l,m,h$ correspond to low, mid, and high exposure settings, respectively, with $m$ serving as the reference image.
-In this work, the default gamma parameter $\gamma$ is set to 2.2.
-Then, we will get a corresponding set of $\left\{H_{i}\right\}$.
-Due to varying luminance levels at different exposure times, captured images may exhibit color rendering deviations from real-world colors, with low-exposure images losing shadow details and high-exposure images losing highlight details.
-Therefore, variations in exposure times certainly affect the overall color accuracy of the input LDR images.
-To address this, we propose a GLA in the pre-alignment subnetwork.
-GLA aims to correct the luminance of LDR images in the sRGB domain, producing consistent brightness features across varying exposures and enhancing detail preservation.
-Specifically, the LDR images are mapped to the luminance range of the reference image and achieve global luminance alignment:
-$G_{i}=clip(L_{i}*(mean(L_{i})/mean(L_{m})),\ i=l,h)$ (2)
-where $G_{i}$ denotes the corresponding luminance-aligned output, $mean(\cdot)$ denotes the mean luminance value, and $clip$ is a function that keeps the output in the range [0, 1].
-Notably, since $L_{m}$ is the reference image, $G_{m}=L_{m}$.
+GLA is introduced to maintain consistent luminance across varying exposures, while the EFSM explores the correlation between inputs and exposure to focus on the most relevant regions for fusion.
+Global Luminance Alignment (GLA).
+Following established HDR reconstruction pipelines [23, 56], we employ three LDR images $\{{L}_{l},{L}_{m},{L}_{h}\}$ (sorted by their exposure length) as input.
+These images are first mapped to the HDR domain relying on gamma correction [23] to produce a set of gamma-corrected images ${G}_{i}$}, where $i\in\{l,m,h\}$ and $\gamma=2.2$.
+To reduce inter-exposure luminance discrepancies and improve alignment robustness, we introduce GLA in the pre-alignment stage.
+GLA operates in the sRGB domain and ensures consistent luminance distributions across inputs and minimizes exposure-induced mismatches during fusion.
+Formally, for each non-reference LDR image $L_{i}$, global alignment is performed as:
+$A_{i}=\operatorname{clip}\!\left(L_{i}\times\frac{\mathbb{E}(L_{m})}{\mathbb{E}(L_{i})},\;0,1\right),\quad i\in\{l,h\},$ (1)
+where $A_{i}$ denotes the luminance-aligned image, $\mathbb{E}(\cdot)$ computes the average luminance, and $\operatorname{clip}(\cdot,0,1)$ restricts pixel values to the valid dynamic range.
+Since the mid-exposure image $L_{m}$ serves as the reference, we set $A_{m}=L_{m}$.
 Figure 4: The pre-alignment subnetwork is composed of two parts: global luminance alignment (GLA) and exposure-guided feature selection module (EFSM).
@@
 As shown in Fig.
-4, given a set of LDR images with varying exposure times, the non-reference images can be aligned to the reference image using GLA.
-Finally, we can obtain 3-pairs of 9-channel tensor $L_{in}=\left(\{L_{l},H_{l},G_{l}\},\{L_{m},H_{m},G_{m}\},\{L_{h},H_{h},G_{h}\}\right)$ as the input to feed into the EFSM module.
-Exposure-Guided Feature Selection Module.
-Existing methods primarily focus on texture for fusion, overlooking the crucial role of exposure time information, which may result in suboptimal fusion and color inconsistencies.
-To overcome this limitation, we propose an EFSM to explore exposure information, enhancing meaningful details and ensuring that the most relevant features are preserved despite varying exposure levels.
-For a set of LDR images with corresponding exposure times of $\left\{{{t_{l}},{t_{m}},{t_{h}}}\right\}$, we standardize and relativize the input exposure times using the following formulation:
-${e_{{}_{l}}}={\log_{2}}\left({{{t_{l}}}/{{t_{m}}}}\right)\cdot c,\ \ \ {e_{{}_{h}}}={\log_{2}}\left({{{t_{h}}}/{{t_{m}}}}\right)\cdot c$ (3)
-where ${e_{{}_{l}}}$ and ${e_{{}_{h}}}$ denote the relative exposure inputs of the low exposure branch and high exposure branch, respectively. $c$ is the hyper-parameter set to 0.1.
-As shown in Fig.
-4, we adopt a multiscale architecture to progressively refine feature representations across different exposure levels.
-The framework initiates with multi-level feature extraction: given input $L_{in}$, three parallel $3\times 3$ convolutional layers generate initial feature maps $\{f_{l},f_{m},f_{h}\}$ corresponding to low-, mid-, and high-exposure features, respectively.
-For subsequent processing clarity, we formally define $f^{nr}\triangleq\{f_{l},f_{h}\}$ as non-reference features and $f^{r}\triangleq f_{m}$ as reference features.
-Assuming there are $N$ scales in total, the reference and non-reference features of the $i$-th scales are denoted as $f_{i}^{r}$ and $f_{i}^{nr}$, respectively ($i=1$ denotes the original scale), $f_{i}^{nr}$ are fed to the shared weights exposure-guided feature selection block (EFSB) and $f_{i}^{r}$ are fed to the shared weights feature selection block (FSB).
-Specifically, the features undergo global average pooling, followed by two fully connected layers and a sigmoid activation to produce the feature-guide modulation coefficients $v^{nr}\in\mathbb{R}^{1\times 1\times 64}$ and $v^{r}\in\mathbb{R}^{1\times 1\times 64}$.
-Simultaneously, relative exposure inputs $e$ are each passed through fully connected layers, resulting in the exposure-guide modulation coefficients $v^{t}\in\mathbb{R}^{1\times 1\times 128}$.
-To dynamically select and refine features from multiple input channels, the exposure-guide modulation coefficients $u^{nr}\in\mathbb{R}^{1\times 1\times 64}$ are computed as follows:
-$u^{nr}=\sigma\left(FC_{3}\left(<v^{nr},v^{t}>\right)\right)$ (4)
-where <,><,>< , > denotes concatenation, $\sigma$ represents sigmoid activation, $FC_{3}$ denotes three fully connected layers. The final output of EFSM can be obtained as:
-$\hat{f_{i}^{r}}=f_{i}^{r}\cdot v^{r},\ \ \ \ \ \ \hat{f_{i}^{nr}}=f_{i}^{nr}\cdot u^{nr}$ (5)
-IV-C Asymmetric Cross-feature Fusion
-While the non-reference image may provide more details, it also poses a higher risk of introducing ghosting artifacts, especially in open scenes.
-To address this, we propose an asymmetric cross-attention block, which enhances the contribution of the reference input during attention computation and mitigates the potential for ghosting artifacts.
-Figure 5: The structure of asymmetric cross-feature block.
-It employs asymmetric cross-attention to align reference and non-reference features, integrating coarse-to-fine guidance for enhanced reference-dominated feature fusion and improved HDR reconstruction.
-Generally, we propose a multi-scale strategy to calculate the attention in a coarse-to-fine manner.
-At each scale, asymmetric cross-attention (ACA) is utilized to compute local attention.
-Specifically, as shown in Fig.
-5, the proposed ACA first utilizes the non-reference feature as an agent, facilitating the exchange of information between the patch tokens of the reference feature and then back project it to its non-reference branch.
-To prioritize mid-exposure image features and encourage more effective exploration of feature interactions across different exposures, we introduce the self-information during the exchange process.
-Following [54], the attention matrix is thus computed as follows:
-$Q_{i}=\phi(\hat{f}_{i}^{r})W_{i}^{q},\ K_{i}=\phi(\hat{f}_{i}^{nr})W_{i}^{k},\ V_{i}=\phi(\hat{f}_{i}^{nr})W_{i}^{v}$ (6)
-$A_{i}=softmax((Q_{i}+K_{i})K_{i}^{T}/\sqrt{C})$ (7)
-where $W_{i}^{q},W_{i}^{k},W_{i}^{v}\in\mathbb{R}^{C\times C}$ are learnable weight matrices of the $i$-th scales, $\phi(\cdot)\in\mathbb{R}^{N\times C}$ denotes the unfolding operation ($N$ refers to the number of patches).
-Then, the aligned non-reference features $f^{anr}_{i}$ are obtained by applying a folding operation to the result of the multiplication of $A_{i}$ and $V_{i}$.
-To leverage attention from coarser scales and local details at the current scale, we introduce the cross-scale guidance feature $f^{guid}$ to achieve coarse-to-fine feature fusion and interaction:
+3, given a set of LDR images with varying exposure times, we construct the input tensor $\mathbf{L}_{\text{in}}=\big{(}\{L_{l},G_{l},A_{l}\},\{L_{m},G_{m},A_{m}\},\{L_{h},G_{h},A_{h}\}\big{)}$, where each triplet $(L_{i},G_{i},A_{i})$ represents the original, gamma-corrected, and GLA-processed versions of the image under exposure setting $i\in\{l,m,h\}$.
+This tensor is then passed to the EFSM for feature selection and alignment refinement.
+Exposure-guided Feature Selection Module (EFSM).
+Exposure information is crucial for distinguishing the reliability and detail across different exposures.
+For the obtained multi-channel features, we propose the EFSM, which consists of the Exposure-guided Feature Selection Block (EFSB) and the Feature Selection Block (FSB).
+EFSM introduces exposure-aware modulation based on relative exposure priors, selectively emphasizing well-exposed regions and suppressing unreliable features, thereby ensuring more reliable feature representation for fusion.
+Specifically, EFSM operates in a multi-scale feature refinement architecture, as depicted in Fig.
+4.
+Given a 9-channel composite input tensor $\mathbf{L}_{\text{in}}$, we first extract three initial feature maps $\{\mathbf{F}_{l},\mathbf{F}_{m},\mathbf{F}_{h}\}$ via parallel $3\times 3$ convolutional layers.
+For subsequent processing clarity, we formally define $\mathbf{F}^{nr}\triangleq\{\mathbf{F}_{l},\mathbf{F}_{h}\}$ as non-reference features and $\mathbf{F}^{r}\triangleq\mathbf{F}_{m}$ as reference features.
+These are fed into two distinct but weight-shared selection blocks: EFSB for non-reference features and FSB for reference features.
+To encode exposure priors, we normalize the exposure times $\{t_{l},t_{m},t_{h}\}$ relative to the middle exposure as:
+$e_{l}=\log_{2}\!\left(\frac{t_{l}}{t_{m}}\right)\cdot c,\quad e_{h}=\log_{2}\!\left(\frac{t_{h}}{t_{m}}\right)\cdot c,$ (2)
+where $e_{l}$ and $e_{h}$ represent the relative exposure embeddings for the low- and high-exposure branches, respectively, and $c$ is a hyper-parameter empirically set to $0.1$.
 […]
-where $f^{guid}_{i}$ represent guidance features at $i$-th scales, $U$ represents pixel shuffle upsampling, $C_{v}$ denotes a $3\times 3$ convolution.
-As shown in Fig.
-5, $f^{guid}_{i}$ learn reference-dominated attention from the previous scale, facilitating the alignment of features across scales.
-The output of asymmetric cross-attention block can be expressed as:
-$\tilde{f}^{nr}=C_{v}(<f^{anr}_{1},U(f^{anr}_{2}),f^{guid}_{1}>)$ (9)
-Finally, we generate a spatial attention map with $\tilde{f}^{nr}$ and $f^{r}$ and get the final output:
-$f^{out}=<\sigma(C_{v}(Relu\left(C_{v}\left(<f^{r},\tilde{f}^{nr}>\right)\right)))\cdot\tilde{f}^{nr},f_{m}>$ (10)
-IV-D Restoration
-The restoration subnetwork adopts a frequency-based approach to effectively mitigate ghosting artifacts in HDR images.
-As shown in Fig.
-3, the subnetwork is designed with a multiscale architecture that progressively refines features at different resolutions.
-The architecture particularly emphasizes feature correction at lower scales, as these scales capture finer high-frequency details, such as textures and small features, which are essential for addressing the residual ghosting artifacts.
-At the lower scales, the features undergo Discrete Wavelet Transform (DWT) to decompose the image into multi-resolution subbands.
-Following this, multiple Learnable Bandpass Filter (LBF) modules [55, 39] are applied to correct each frequency component.
-The output is then passed through the Inverse Wavelet Transform (IWT) to reconstruct the feature maps $f_{\times 1}$ in the feature domain.
-At higher scales, after applying DWT, a $1\times 1$ convolution is used instead of multiple LBF modules.
-The feature maps at higher scales primarily contain global structural information, reducing the need for complex high-frequency corrections.
-Finally, we combine the high-frequency details from lower scales with the global information from higher scales, leading to improved HDR image restoration with reduced ghosting artifacts:
-$H_{out}=\sigma(Conv(U(U(f_{\times 4})+f_{\times 2})+f_{\times 1}))$ (11)
-where $U$ denotes upsampling, $f_{\times 4},f_{\times 2},f_{\times 1}$ are the output features at different scales, $H_{out}$ represents the final HDR image output.
+$\mathbf{u}^{\text{nr}}=\sigma\!\big{(}\operatorname{FC}_{3}([\mathbf{v}^{\text{nr}};\mathbf{v}^{\text{t}}])\big{)},$ (3)
+where $\cdot$$\cdot$] denotes channel-wise concatenation, $\sigma(\cdot)$ is sigmoid activation, and $\operatorname{FC}_{3}(\cdot)$ represents three fully connected layers.
+Finally, the refined features are obtained by reweighting:
+$\hat{\mathbf{F}}_{i}^{\text{r}}=\mathbf{F}_{i}^{\text{r}}\odot\mathbf{v}^{\text{r}},\quad\hat{\mathbf{F}}_{i}^{\text{nr}}=\mathbf{F}_{i}^{\text{nr}}\odot\mathbf{u}^{\text{nr}},$ (4)
+where $\odot$ denotes element-wise multiplication.
+IV-C asymmetric cross-feature fusion subnetwork
+One of the central challenges in multi-exposure fusion lies in leveraging the complementary details from non-reference images without introducing ghosting artifacts, particularly in the presence of large motion or occlusions [58, 34].
+While non-reference inputs offer richer information in under-/over-exposed regions, they also carry a higher risk of misalignment.
+To address this, we propose an asymmetric cross-feature fusion subnetwork, which employs asymmetric cross-feature fusion blocks to selectively align and fuse cross-exposure features in a coarse-to-fine manner.
+Each fusion block is centered around the Asymmetric Cross-Attention (ACA) module, which leverages reference-dominant attention to fuse features.
+Figure 5: The structure of the asymmetric cross-feature fusion block.
+It employs asymmetric cross-feature attention to align reference and non-reference features, integrating coarse-to-fine guidance for enhanced reference-dominated feature fusion and improved HDR reconstruction.
+As illustrated in Fig.
+5, the proposed module adopts a multi-scale architecture, where local attention at each level is computed via an ACA mechanism.
+At each scale $i\in\{1,2,3\}$, we first unfold the refined reference features $\hat{\mathbf{F}}_{i}^{\text{r}}$ and non-reference features $\hat{\mathbf{F}}_{i}^{\text{nr}}$ into patch tokens using a window-based unfolding operator $\phi(\cdot)$.
+The attention is formulated in an asymmetric manner, where queries are derived from the reference stream, while keys and values are obtained from the non-reference stream to facilitate reference-guided feature aggregation.
+To further emphasize reference dominance, we incorporate query features into the key projection via additive modulation:
+$\mathbf{M}_{i}=\operatorname{softmax}\!\left((\mathbf{Q}_{i}+\mathbf{K}_{i})\mathbf{K}_{i}^{\top}/{\sqrt{C}}\right),$ (5)
+where $\mathbf{Q}_{i}=\phi(\hat{\mathbf{F}}_{i}^{\text{r}})\,\mathbf{W}_{i}^{q}$ and $\mathbf{K}_{i},\mathbf{V}_{i}=\phi(\hat{\mathbf{F}}_{i}^{\text{nr}})\,\mathbf{W}_{i}^{k,v}$, with $\mathbf{W}_{i}^{q,k,v}$ being learnable matrices.
+This asymmetric formulation encourages stronger alignment toward the reference features and reduces the impact of unreliable non-reference content.
+Then, the aligned non-reference features $\mathbf{F}^{anr}_{i}$ are obtained by applying a folding operation to the result of the multiplication of $\mathbf{M}_{i}$ and $\mathbf{V}_{i}$.
+To enhance fine-scale alignment with global context, we incorporate a cross-scale guidance mechanism.
+Specifically, attention information from the coarser scale $i+1$ is projected and injected into the finer scale $i$ as:
+$\mathbf{F}_{i}^{\text{guid}}=\operatorname{Conv}_{3\times 3}\big{(}[\mathbf{F}_{i}^{\text{nr}};U(\psi(\mathbf{M}_{i+1}))]\big{)},\quad 1\leq i<N,$ (6)
+where $\operatorname{Conv}_{3\times 3}(\cdot)$ denotes a $3\times 3$ convolution, $N$ denotes the number of refinement scales, $U(\cdot)$ is pixel shuffle upsampling, and $\psi(\cdot)$ represents the projection of attention from the coarser scale.
+Then, $\mathbf{F}_{i}^{\text{guid}}$ injects reference-dominated attention to facilitate multi-scale alignment.
+The output of the ACA $\mathbf{F}^{\text{nr}}$ is computed by merging the aligned features from multiple scales, combining the finer-scale features with guidance from the coarser scales to produce a refined feature representation.
+Finally, we generate a spatial attention map using $\tilde{\mathbf{F}}^{\text{nr}}$ and $\mathbf{F}^{\text{ref}}$ and get the fusion output.
+IV-D restoration subnetwork
+As shown in Fig. 3, the restoration subnetwork is designed to suppress ghosting artifacts while preserving fine details across multiple scales.
+Each scale begins with a Discrete Wavelet Transform (DWT) to decompose features into frequency subbands.
+At the lowest scale, where high-frequency details such as textures and residual motion artifacts are most prominent, we apply a Lightweight DomainPlus Block (LDPB) to perform frequency-specific corrections.
+The LDPB is adapted from [62] by removing dense connections to reduce complexity while retaining frequency modeling capacity.
+In contrast, higher scales capture more global and low-frequency content.
+These layers are processed using a simple $1\times 1$ convolution after DWT, which is sufficient for coarse structure refinement without the need for full spectral filtering.
+All features are subsequently reconstructed using the Inverse Wavelet Transform (IWT).
+The refined features from each scale are fused in a top-down manner using upsampling and residual connection to produce the final HDR output.
 IV-E Loss Function
@@
 To enhance training performance, we apply the $\mu$-law transformation, as proposed in [22], to map the HDR image from the linear domain to the tone-mapped domain, thereby improving model performance during training, the $\mu$-law transformation can be expressed as:
-$\tau(H)=\frac{\ln(1+\mu H)}{\ln(1+\mu)}$ (12)
+$\tau(\mathcal{H})=\frac{\ln(1+\mu\mathcal{H})}{\ln(1+\mu)}$ (7)
 where $\mu$ is a hyperparameter set to 5000.
@@
 To correct each frequency component and enhance the training of the restoration subnetwork, we adopt the dilated advanced Sobel loss (D-ASL) loss function [63] as an additional function, which is defined as:
-$\mathcal{A}\mathcal{S}\mathcal{L}(\widehat{Z},Z)=\frac{1}{N}\sum\mid\operatorname{Sobel}^{\star}(\widehat{Z})-\text{ Sobel }^{\star}(Z)\mid\$ (13)
+$\mathcal{A}\mathcal{S}\mathcal{L}(\widehat{Z},Z)=\frac{1}{N}\sum\mid\operatorname{Sobel}^{\star}(\widehat{Z})-\text{Sobel}^{\star}(Z)\mid\\$ (8)
 ${\cal D}-{\cal A}{\cal S}{{\cal L}^{\left\{{{d_{1}},{d_{2}},\cdots,{d_{n}}}\right\}}}=\sum\limits_{i=1}^{n}{{\cal A}{\cal S}{\cal L}\left|{{}_{dilation\_rate={d_{i}}}}\right.}$ (9)
@@
 Then the final loss function can be expressed as:
-$\operatorname{Loss}(\widehat{Z},Z)=\mathcal{L}1(\widehat{Z},Z)+\lambda\cdot\mathcal{D}-\mathcal{AS}\mathcal{L}^{D}(\widehat{Z},Z)$ (15)
+$\mathcal{L}_{total}=\mathcal{L}1(\widehat{Z},Z)+\lambda\cdot\mathcal{D}-\mathcal{AS}\mathcal{L}^{D}(\widehat{Z},Z)$ (10)
 where $\lambda$ = 0.25 is a hyper-parameter to balance the L1 loss and D-ASL.
+TABLE I: Comparison results on Kalantari’s dataset and Prabhakar’s dataset.
+Throughout this paper, the best and second-best results of each case are highlighted in bold red and underlined blue, respectively.
+dataset setting train and test on Kalantari’s dataset train and test on Prabhakar’s dataset
+Method metrics PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$) HDR-VDP-2($\uparrow$) PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$) HDR-VDP-2 ($\uparrow$)
+Kalantari(CGF 2017) [23] 42.74 41.22 0.9877 0.9848 60.51 35.63 32.50 0.09613 0.9692 59.42
+AHDRNet(CVPR 2019) [56] 43.77 41.35 0.9907 0.9859 62.30 38.61 35.26 0.9663 0.9794 61.14
+Prabhakar(ECCV 2020) [43] 43.08 41.68 - - 62.21 38.30 34.98 0.9702 0.9781 -
+HDR-Trans(ECCV 2022) [34] 44.28 42.88 0.9916 0.9884 66.03 41.31 39.44 0.9726 0.9885 63.01
+DomainPlus(MM 2022) [62] 44.02 41.28 0.9910 0.9864 62.91 40.38 38.08 0.9698 0.9872 62.12
+SCTNet(ICCV 2023) [46] 44.13 42.12 0.9916 0.9890 66.65 41.23 38.75 0.9724 0.9881 62.29
+SAFNet(ECCV 2024) [27] 44.61 43.09 0.9918 0.9892 66.93 40.18 37.90 0.9705 0.9865 62.04
+EAFNet(Ours) 44.69 42.19 0.9920 0.9895 68.35 41.80 40.13 0.9731 0.9895 63.53
+TABLE II: Cross-dataset evaluation on Kalantari’s dataset and Prabhakar’s dataset.
+dataset setting train on Kalantari’s dataset, test on Prabhakar’s dataset train on Prabhakar’s dataset, test on Kalantari’s dataset
+Method metrics PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$) PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$)
+AHDRNet(CVPR 2019) [56] 33.96 32.46 0.9601 0.9542 40.03 36.71 0.9855 0.9758
+HDR-Trans(ECCV 2022) [34] 34.07 36.62 0.9675 0.9656 41.38 39.21 0.9890 0.9873
+DomainPlus(MM 2022) [62] 32.64 30.42 0.9046 0.9074 41.15 38.18 0.9873 0.9837
+SCTNet(ICCV 2023) [46] 33.83 30.95 0.9584 0.9521 40.88 37.59 0.9892 0.9842
+SAFNet(ECCV 2024) [27] 38.00 34.65 0.9597 0.9793 40.86 37.50 0.9882 0.9810
+EAFNet(Ours) 39.26 35.99 0.9707 0.9848 42.02 39.38 0.9903 0.9870
 V EXPERIMENTS
-V-1 Settings
+In this section, we conduct extensive experiments to evaluate the performance of our proposed dual-stream paradigm in HDR video generation.
+Our dual-stream paradigm, supported by a dual-camera system and the EAFNet model, aims to address key challenges in HDR video reconstruction, such as temporal luminance consistency and spatial deghosting.
+First, in Section V-A, we outline the experimental setup, including details on the training process, datasets, baseline methods, and evaluation metrics.
+In Section V-B, we focus on dynamic HDR image fusion, presenting both quantitative and qualitative comparisons with state-of-the-art HDR image deghosting methods to validate the effectiveness of our method in single-frame spatial fusion and deghosting.
+Section V-C shifts to HDR video reconstruction, where we assess the advantages of our paradigm in temporal consistency and flicker suppression, comparing it with existing AE paradigms and further validating the benefits of the EAFNet model within our proposed paradigm.
+Section V-D presents an ablation study, analyzing the impact of the dual-camera system configuration and the contributions of each EAFNet component.
+Finally, Section V-E discusses the limitations of the proposed paradigm, including the effects of camera parallax and system efficiency, and highlights potential directions for future improvements.
+V-A Experiment Settings
+Training Details.
 The proposed method is implemented in PyTorch on an NVIDIA RTX4090 GPU.
-For the convolutional layers, we employ 64 kernels of size 3$\times$3 with a stride of one, applying zero padding to preserve the dimensions of the resulting feature maps.
+For the convolutional layers, we employ 64 kernels of size $\times3$ with a stride of one, applying zero padding to preserve the dimensions of the resulting feature maps.
 The optimization is performed using the Adam optimizer [26] with an initial learning rate of $10^{-4}$, and the training concludes when the learning rate reaches $10^{-6}$.
@@
 Patches are randomly rotated for data augmentation to avoid overfitting during the training stage.
-V-2 Datasets
-We employ four benchmark datasets for evaluation: Prabhakar’s dataset [30], Kalantari’s dataset [28], Tursun’s dataset [58] and Sen’s dataset [5]. Details of these datasets are described as follows:
-1). Prabhakar’s Dataset. Including 582 varying exposure images with corresponding deghosted HDR images, with exposure values between -3EV to +3EV.
-2). Kalantari’s dataset. Including 74 training samples and 15 testing samples. Each consists of three LDR images, and the exposure values belong to {-2, 0, +2} or {-3, 0, +3}.
-3). Tursun’s dataset and Sen’s dataset. Lacking ground truth, is used to assess the generalization ability of the models.
-V-3 Metrics
+Datasets.
+High-quality HDR video datasets suitable for supervised training or frame-wise quantitative evaluation are scarce, as most lack per-frame HDR ground truth and sufficient exposure diversity.
+To fairly and comprehensively evaluate our method, we separate the evaluation into two complementary parts: (1) Spatial evaluation focuses on single-frame fusion accuracy using HDR image datasets with reliable ground truth, enabling objective comparison in reconstruction quality.
+We assess single-frame fusion quality using three widely used HDR image datasets: Prabhakar’s dataset [44], Kalantari’s dataset [22], and Tursun’s dataset [48],
+  • Prabhakar’s Dataset. Including 466 training samples and 116 testing samples, with exposure values between -3EV and +3EV.
+  • Kalantari’s dataset. Including 74 training samples and 15 testing samples. Each consists of three LDR images, with exposure values of {-2 EV, 0 EV, +2 EV} or {-3 EV, 0 EV, +3 EV}.
+  • Tursun’s dataset. Lacking ground truth, is used to assess the generalization ability of the models.
+(2) Temporal evaluation focuses on luminance consistency and flicker suppression in dynamic scenes, which cannot be assessed using static image datasets.
+To enable a comprehensive comparison of our dual-stream HDR video generation paradigm against the widely adopted AE paradigm, we construct dedicated datasets tailored to each paradigm under real-world scenes using our DCS.
+This approach is necessary because the two paradigms differ fundamentally in their acquisition strategies; a direct comparison using identical inputs would be inherently unfair.
+Specifically, we capture two synchronized video streams for each scene using two identical cameras (resolution $960\times 540$).
+This synchronization is applied only during data acquisition to minimize scene variations between paradigms.
+In practical deployment, our dual-camera system operates asynchronously without hardware-level synchronization (see Section III).
+The first camera follows an alternating exposure pattern, and the second camera continuously captures medium-exposure frames.
+Our paradigm uses the low- and high-exposure frames from the first camera together with the medium-exposure frames from the second camera to reconstruct HDR video.
+For the AE pipeline, only the alternating exposure sequence from the first camera is used, emulating the standard AE pipeline.
+We capture 26 real-world scenes under matched conditions for both datasets, covering diverse scenarios including indoor/outdoor environments, day/night lighting, and challenging situations such as low-light, high-contrast scenes, object motion, and occlusions.
+Exposure settings are fixed to (2,0,+2)(-2,0,+2) or (1,0,+1)(-1,0,+1) EV to ensure sufficient exposure diversity across varying lighting conditions.
+This paired capture protocol guarantees that both paradigms are evaluated on scene-matched content while preserving their respective optimal acquisition strategies.
+However, since the dataset is captured in real-world scenes, there are two challenges in ensuring a fair comparison.
+First, slight parallax exists between the two camera streams, as the cameras are not perfectly aligned.
+To mitigate this, we capture predominantly long-range scenes, which minimize the impact of parallax.
+Second, temporal asynchrony arises from the exposure duration differences inherent in the AE settings.
+This asynchrony is a fundamental aspect of the AE paradigm and cannot be avoided.
+With these considerations, our paired capture protocol ensures that both paradigms are evaluated on scene-matched content while preserving their respective optimal acquisition strategies.
+Figure 6: Visual comparison on Kalantari’s and Prabhakar’s dataset.
+While other methods suffer from detail loss, overexposed patches, and color inconsistencies, our approach successfully reconstructs overexposed details and preserves natural, consistent color.
+Figure 7: Visual comparison on Tursun’s dataset.
+Existing methods often produce artifacts due to luminance contrast between reference and non-reference images, while SAFNet effectively suppresses these artifacts.
+Baselines.
+We quantitatively and qualitatively compare our EAFNet to several state-of-the-art HDR image deghosting methods including Kalantari [23], AHDR [56], Prabhakar [43], HDR-Transformer [34], SCTNet [46], DomainPlus [62], and SAFNet [27].
+For fairness, we exclude few-shot and unsupervised methods that are not designed for supervised multi-exposure HDR reconstruction.
+We further compare our proposed framework with representative AE-based HDR video reconstruction methods, including HDRFlow [54], LAN-HDR [8], and DeepHDRVideo [2].
+Due to differences in system assumptions and training strategies, AE methods are trained on their respective public datasets following the official configurations, while our method is trained on the publicly available Prabhakar dataset.
+All models are evaluated on the same real-world video sequences captured using our dual-camera system.
+Metrics.
 We quantitatively evaluate the predicted HDR images by measuring PSNR and SSIM [52] in both the linear domain (-L) and the tone-mapped domain (-$\mu$) [22].
@@
     where $L_{i}$ is the mean luminance of frame $i$.
-V-A Comparison with state-of-the-art methods
-We quantitatively and qualitatively compare our EAFNet to several state-of-the-art MEF methods including Kalantari [49], AHDR [36], Prabhakar [61], HDR-Transformer [40], SCTNet [43], DomainPlus [39], SAFNet [31], and several state-of-the-art AE methods including HDRFlow [51], LAN-HDR [50], and DeepHDRVideo [19].
-TABLE I: Comparison results on Kalantari’s dataset and Prabhakar’s dataset. The best results are highlighted and the second best results are underlined.
-Models 17’CGF 19’CVPR 20’ECCV 22’ECCV 22’MM 23’ICCV 24’ECCV
-Kalantari AHDRNet Prabhakar HDR-Trans DomainPlus SCTNet SAFNet Ours
-train and test on Kalantari’s dataset
-PSNR-µ 42.74 43.77 43.08 44.28 44.02 44.13 44.61 44.69
-PSNR-L 41.22 41.35 41.68 42.88 41.28 42.12 43.09 42.19
-SSIM-µ 0.9877 0.9907 / 0.9916 0.9910 0.9916 0.9918 0.9920
-SSIM-L 0.9848 0.9859 / 0.9884 0.9864 0.9890 0.9892 0.9895
-HDR-VDP-2 60.51 62.30 62.21 66.03 62.91 66.65 66.93 68.35
-train and test on Prabhakar’s dataset
-PSNR-µ 35.63 38.61 38.30 41.31 40.38 41.23 40.18 41.80
-PSNR-L 32.50 35.26 34.98 39.44 38.08 38.75 37.90 40.13
-SSIM-µ 0.9613 0.9663 0.9702 0.9726 0.9698 0.9724 0.9705 0.9731
-SSIM-L 0.9692 0.9794 0.9781 0.9885 0.9872 0.9881 0.9865 0.9895
-HDR-VDP-2 59.42 61.14 / 63.01 62.12 62.29 62.04 63.53
-TABLE II: Cross-dataset evaluation on Kalantari’s dataset and Prabhakar’s dataset. The best results are highlighted and the second best results are underlined.
-2019 2022 2022 2023 2024
-Models AHDRNet HDR-Trans DomainPlus SCTNet SAFNet Ours
-train on Kalantari’s dataset, test on Prabhakar’s dataset
-PSNR-µ 33.96 34.07 32.64 33.83 38.00 39.26
-PSNR-L 32.46 36.62 30.4233 30.95 34.65 35.99
-SSIM-µ 0.9601 0.9675 0.9046 0.9584 0.9597 0.9707
-SSIM-L 0.9542 0.9656 0.9094 0.9521 0.9793 0.9848
-train on Prabhakar’s dataset, test on Kalantari’s dataset
-PSNR-µ 40.03 41.38 41.15 40.88 40.86 42.02
-PSNR-L 36.71 39.21 38.18 37.59 37.50 39.38
-SSIM-µ 0.9855 0.9890 0.9873 0.9892 0.9882 0.9903
-SSIM-L 0.9758 0.9873 0.9837 0.9842 0.9810 0.9870
-Figure 6: Visual comparison on Kalantari’s dataset.
-While other methods suffer from detail loss, overexposed patches, and color inconsistencies, our approach successfully reconstructs overexposed details and preserves natural, consistent color.
-Figure 7: Visual comparison on Prabhakar’s dataset. Most existing methods suffer from ghosting artifacts, our EAFNet mitigates these issues.
-Experiments on Prabhakar’s datasets and Kalantari’s datasets.
-Intra-Validation.
-Our DCS can be viewed as a MEF task, so we conduct a comparison with the state-of-the-art MEF method using the publicly available MEF dataset.
-The quantitative results on both datasets are shown in Table I, our EAFNet achieves superior performance across all datasets.
-Specifically, on the Kalantari dataset, our method surpasses the second-best approach in terms of PSNR-$\mu$ by 0.08 dB, while on the Prabhakar dataset, it achieves a more substantial improvement of 0.49 dB.
-The visual comparisons in Fig.
-7 illustrate that most existing methods still suffer from noticeable ghosting artifacts.
-In contrast, the proposed EAFNet effectively mitigates these issues.
-Moreover, in overexposed regions (Figure 6), we find that the Kalantari method leads to detail loss, while AHDRNet and HDR-Trans exhibit large white overexposed patches in wall areas.
-Additionally, methods such as DomainPlus still struggle with color inconsistencies.
-In contrast, our approach successfully reconstructs most of the overexposed details while preserving natural and consistent color reproduction.
-Figure 8: Visual comparison on Tursun’s dataset.
-Existing methods often produce artifacts due to luminance contrast between reference and non-reference images, while SAFNet effectively suppresses these artifacts.
-Figure 9: Visual comparison on Sen’s dataset.
-Existing methods suffer from misalignment or detail loss when the reference image is over-exposed, while our method reconstructs clearer and more accurate textures.
-Cross-Validation.
-The cross-testing on the two datasets serves to mitigate any potential bias and offers a comprehensive understanding of the model’s reliability across varying conditions.
-As shown in Table II, we achieve superior performance in cross-dataset evaluation, demonstrating the robustness and generalizability of our model.
-Unsupervised-dataset.
-To assess the generalization performance of our SAFNet, we conduct qualitative comparisons on the Sen’s and Tursun’s datasets.
-As shown in Fig.
-8, when there exists a significant luminance contrast between the reference and non-reference images, such as the case where white backgrounds in the reference image correspond to dark content in the non-reference frames, most existing methods tend to produce prominent visual artifacts.
-In contrast, the proposed SAFNet effectively suppresses artifacts, delivering perceptually plausible HDR reconstructions even in scenes with severe exposure disparity.
-In Fig.
-9, When the reference image is over-exposed and misalignment exists between the reference and non-reference images, methods such as AHDRNet, DominPlus, and SAFNet exhibit incorrect alignment, while HDR-Trans and SCTNet suffer from varying degrees of detail loss.
-In contrast, our method is capable of reconstructing clearer and more accurate textures.
-Figure 10: Visual comparison with MEF methods on self-captured videos.
+TABLE III: Comparison on self-captured videos.
+* indicates HDR image deghosting methods under the alternating exposure solution.
+The reported time refers to the inference time for a 128×128 patch on a single RTX 4090 GPU.
+Method Paradim LSD ($\downarrow$) t-SSIM ($\uparrow$) $L_{avg}$ Inference time (ms, $\downarrow$)
+DeepHDRVideo(ICCV 2021) [2] 0.1031 0.5913 105.02 22.23
+LAN-HDR(ICCV 2023) [8] 0.1003 0.6879 108.75 21.69
+HDRFlow(CVPR 2024) [54] 0.0914 0.7352 113.49 2.62
+SCTNet*(ICCV 2023) [46] 0.1027 0.6947 109.23 11.17
+SAFNet*(ECCV 2024) [27] alternating-exposure paradigm 0.0976 0.6862 111.70 3.27
+HDR-Trans(ECCV 2022) [34] 0.0093 0.8923 111.24 21.60
+DomainPlus(MM 2022) [62] 0.0094 0.8991 110.41 7.44
+SCTNet(ICCV 2023) [46] 0.0093 0.8868 106.84 11.17
+SAFNet(ECCV 2024) [27] our dual-stream paradigm 0.0097 0.8851 111.32 3.24
+EAFNet(Ours) 0.0093 0.9071 133.16 4.32
+V-B HDR Image Experimental Results
+Quantitative Comparisons.
+Table I reports the intra-dataset evaluation results on Prabhakar’s and Kalantari’s datasets.
+Our EAFNet consistently achieves the best performance across both datasets.
+On the Kalantari dataset, it surpasses the second-best method by $0.08$ dB in PSNR-$\mu$, while on the Prabhakar dataset, the margin increases to $0.49$ dB.
+These gains are complemented by improvements in SSIM-$\mu$, indicating that our exposure-adaptive fusion strategy benefits both fidelity and structural consistency.
+We further conduct cross-dataset validation, where training and testing are performed on different datasets (Table II).
+This setting is critical for evaluating the robustness of HDR fusion models against variations in exposure patterns, motion distributions, and scene content, thereby better reflecting real-world deployment scenarios.
+Our EAFNet maintains clear superiority in this challenging setting, with cross-domain gains even larger than in the intra-dataset case.
+For example, when trained on Kalantari and tested on Prabhakar, EAFNet achieves 39.26 dB PSNR-$\mu$ and 0.9707 SSIM-$\mu$, surpassing the next best approach by over 1.2 dB and 0.03, respectively.
+This demonstrates that our model does not overfit to dataset-specific statistics, but learns exposure-aware and motion-robust fusion representations that transfer effectively across domains.
+The strong bidirectional results confirm the generality and domain-agnostic nature of our fusion mechanism.
+Qualitative Comparisons.
+Fig.6 presents visual comparisons on the Prabhakar and Kalantari datasets.
+Most existing methods exhibit noticeable ghosting artifacts, particularly around moving objects and edges with exposure differences.
+In overexposed regions, the Kalantari method suffers from detail loss, while AHDRNet and HDR-Trans produce large saturated patches, and DomainPlus struggles with color inconsistencies.
+In contrast, EAFNet recovers fine structures in saturated areas while preserving natural and consistent color reproduction.
+We also evaluate on challenging HDR datasets without ground truth.
+In Tursun’s dataset (Fig.7), large luminance differences between reference and non-reference frames often cause severe artifacts in competing methods, particularly when bright backgrounds in the reference correspond to dark content in non-reference frames.
+EAFNet effectively suppresses these artifacts, producing perceptually coherent HDR reconstructions.
+Figure 8: Visual comparison with HDR image deghosting methods on self-captured videos.
 Our EAFNet reduces ghosting from motion (first row) and background misalignment (second row), recovering natural colors and fine details in dark regions (third row).
 Figure 9: Visual comparison with AE methods on self-captured video dataset. Our dual-camera system reconstructs HDR video with consistent luminance.
-TABLE III: Comparison on self-captured videos. * indicates MEF methods under the AE setup. The reported time refers to the inference time for a 128×128 patch on a single RTX 4090 GPU.
-AE setup MEF setup with DCS
-21’ICCV 23’ICCV 24’CVPR 23’ICCV 24’ECCV 22’ECCV 22’MM 23’ICCV 24’ECCV
-Chen LAN-HDR HDRFlow SCTNet* SAFNet* HDRTrans DomainPlus SCTNet SAFNet Ours
-LSD 0.1031 0.1003 0.0914 0.1027 0.0976 0.0093 0.0094 0.0093 0.0097 0.0093
-t-SSIM 0.5913 0.6879 0.7352 0.6947 0.6862 0.8923 0.8991 0.8868 0.8851 0.9071
-$L_{avg}$ 105.02 108.75 113.49 109.23 111.70 111.24 110.41 106.84 111.32 133.16
-Runtime(ms) 22.23 21.69 2.6 11.17 3.2 21.6 7.44 11.17 3.2 4.3
-V-B Experiments on self-captured videos
-Overview.
-In this section, we explore the effectiveness of our DCS in real-world scenes and assess the capabilities of the proposed EAFNet in HDR video reconstruction.
-As detailed in Section III, real video data is captured using dual MV-CS032-10GC industrial cameras.
-Due to the alternating exposure method’s requirement for continuous alternating exposure frames, direct comparison with our approach under identical settings is not feasible.
-To ensure a fair comparison, we simultaneously capture two datasets with our system, ensuring minimal scene variation.
-The first camera captures alternating low, medium, and high exposure sequences, while the second camera continuously captures a medium exposure sequence.
-The AE method uses the alternating exposure sequence from the first camera as the test dataset, while our dual-camera framework takes low and high-exposure images from the first camera and a continuous medium-exposure sequence from the second.
-Both cameras have a resolution of $960\times 540$, and we select nine scenes spanning indoor, outdoor, day, and night conditions, each covering low-light, normal-light, and high-light scenarios with challenges such as high contrast, motion, and occlusion.
-The gain follows the scene variations, using exposure times of (-2, 0, 2 EV).
-Compare with MEF methods.
-We compare our method with the state-of-the-art MEF methods including HDR-Transformer [40], SCTNet [43], DomainPlus [39], and SAFNet [31].
-Firstly, we compare the performance of models trained on two different datasets, Prabhakar’s and Kalantari’s.
+V-C HDR Video Experimental Results
+Quantitative Comparisons. Table III summarizes the quantitative results. We interpret these findings from two perspectives:
+Paradigm-level comparison (AE vs.
+our dual-stream).
+Dual-stream-based methods consistently achieve lower luminance standard deviation (LSD) and higher temporal SSIM (t-SSIM) than AE-based methods, indicating superior temporal stability and reduced flickering.
+For instance, the best AE-based method, HDRFlow, attains an LSD of 0.0914 and a t-SSIM of 0.7352, whereas our EAFNet under the dual-stream paradigm achieves 0.0093 and 0.9071, respectively.
+This substantial improvement stems from the fixed-exposure temporal reference in the dual-stream pipeline, which effectively eliminates inter-frame luminance flickering caused by alternating exposure in AE pipelines.
+Within-dual-stream comparison (our EAFNet vs.
+image deghosting baselines).
+When all methods operate under our dual-stream paradigm, our EAFNet outperforms all integrated image deghosting-based baselines in both temporal and spatial quality metrics.
+EAFNet achieves the highest t-SSIM (0.9071) and ties for the best LSD (0.0093), reflecting its ability to simultaneously preserve temporal coherence and fine-grained spatial details.
+Notably, it accomplishes this while maintaining a reasonable inference time (4.32 ms for a $128\times 128$ patch), which is faster than most transformer-based baselines.
+Qualitative Comparisons. Visual comparisons in Fig. 9 and Fig. 8 further support the quantitative findings.
+Paradigm-level comparison (AE vs.
+our dual-stream).
 As shown in Fig.
-10 (third row), we find that models trained on Kalantari’s dataset exhibit unnatural textures in certain scenes.
-Therefore, all MEF results are tested using the model trained on Prabhakar’s dataset.
-During real video capture, complex lighting conditions, camera and object motion, and high-contrast scenes pose significant challenges for video reconstruction.
-As shown in the first row of Fig.
-10, DomainPlus produces motion misalignment artifacts, while SAFNet and STCNet generate unnatural artifacts in some very dark regions.
-In contrast, our EAFNet effectively eliminates these artifacts, restoring more natural colors and brightness in darker regions, thus providing more detailed information.
-The second row of the figure demonstrates that for high-contrast scenes (where reference images are overexposed or areas with white backgrounds correspond to low/high exposure images of dark objects), existing methods fail to eliminate misalignment artifacts.
-However, our EAFNet can produce more natural reconstruction results, delivering a visually coherent output.
-The third row shows that our EAFNet provides more detail in the dark region, which leads to high luminance in Table III.
-Compare with AE methods.
-Since the AE method and our method require training on different dataset settings, the AE method is trained on its public dataset following the AE training setup, while our method is trained on the publicly available Prabhakar’s dataset.
-The models are then tested on our dual-camera dataset.
-We compare the state-of-the-art AE methods including HDRFlow [51], LAN-HDR [50], and DeepHDRVideo [19].
+9, AE-based reconstructions exhibit noticeable luminance fluctuations between consecutive frames.
+These artifacts are inherent to AE pipelines due to the lack of a consistent exposure reference.
+In contrast, our results maintain stable luminance across frames and significantly reduce temporal artifacts, benefiting from the persistent medium-exposure reference stream.
+Within-dual-stream comparison (our EAFNet vs.
+image deghosting baselines).
 As shown in Fig.
-11 and Table III, for real-world test videos captured in our DCS, the luminance of the video reconstructed using the AE method exhibits abrupt changes, leading to noticeable flickering, which makes it impractical for real-world applications.
-In contrast, our DCS produces temporally consistent results, which suggests that the proposed system can effectively overcome the challenges of traditional AE methods.
-Efficiency Analysis.
-The overall efficiency of HDR video processing is determined by both the acquisition throughput and the inference speed.
-Our method specifically targets the acquisition bottleneck by introducing an asynchronous dual-camera design, which decouples the capture of alternate exposure frames from the reference stream.
-This architecture effectively eliminates the latency caused by long high-exposure frames in traditional single-camera systems, enabling real-time HDR frame acquisition.
-Although our reconstruction model (EAFNet) is not specifically optimized for real-time inference, as shown in Table III, it still exhibits competitive runtime performance.
-Dual-Camera Analysis.
-The use of a dual-camera system inevitably introduces parallax and temporal misalignment due to the baseline separation and potential desynchronization between the two cameras.
-In our prototype, the baseline is constrained by mechanical limitations but is minimized to the extent possible.
-Although some parallax is unavoidable, we find that most MEF-based methods in our system demonstrate reasonable robustness to mild parallax.
-The training data involves dynamic scenes with moving objects and occlusions, the model thus learns to handle motion-induced spatial misalignments, which share similar characteristics with parallax-related shifts in dual-camera setups.
-In challenging cases, such as those involving high local contrast, misalignment becomes more difficult to resolve.
-For instance, in the second row of Figure 10, the presence of strong contrast across exposure frames poses a severe test for existing methods.
-In typical dynamic scenes with such conditions, most existing approaches struggle to produce satisfactory results.
-In contrast, our proposed EAFNet effectively mitigates artifacts, preserving detail and structure even under these difficult situations.
-This robustness is achieved through a combination of a pre-alignment module and a reference-guided fusion strategy, which together reduce visible artifacts even in scenes with high contrast or occlusion.
-In addition, the system does not rely on precise camera calibration or strict synchronization.
-The fusion framework is designed to tolerate both geometric misalignment and temporal offset.
-As shown in Table IV, ablation studies under various calibration conditions and randomized non-reference input orders confirm that the overall reconstruction quality remains stable.
-Small changes in t-SSIM are mainly due to image cropping rather than alignment failure.
-TABLE IV: t-SSIM comparison under different dual-camera deployment settings
-t-SSIM MEF setup with DCS
-22’ECCV 22’MM 23’ICCV 24’ECCV
-HDRTrans DomainPlus SCTNet SAFNet Ours
-ori 0.8923 0.8991 0.8868 0.8851 0.9071
-w/o calibration 0.8922 0.8994 0.8872 0.8861 0.9079
-random non-reference 0.8926 0.8997 0.8842 0.8848 0.9072
-lower frame rate 0.8924 0.8985 0.8871 0.8836 0.9066
-TABLE V: The ablation study of pre-alignment subnetwork and asymmetric cross-feature fusion subnetwork on Prabhakar’s datasets
-pre align fusion PSNR-$\mu$ PSNR-L SSIM-$\mu$ SSIM-L
-GLA EFSM ACA $f^{guide}$
-$\surd$ $\surd$ 41.29 39.47 0.9728 0.9888
-$\surd$ $\surd$ $\surd$ 41.43 39.73 0.9722 0.9890
-$\surd$ $\surd$ $\surd$ 41.17 39.33 0.9726 0.9780
-$\surd$ $\surd$ $\surd$ $\surd$ 41.80 40.14 0.9731 0.9895
-$\surd$ $\surd$ CA $\surd$ 41.61 39.80 0.9726 0.9892
-$\surd$ $\surd$ $\surd$ 41.48 39.79 0.9723 0.9893
-$\surd$ $\surd$ 41.13 39.12 0.9708 0.9887
-40.81 38.70 0.9703 0.9784
+8, in high-motion scenarios, some image deghosting baselines (e.g., DomainPlus, SAFNet) generate ghosting or introduce color distortions in dark regions, whereas our EAFNet preserves spatial alignment and reproduces accurate color.
+Under extreme luminance contrast, EAFNet better suppresses misalignment artifacts, maintaining natural transitions in overexposed or underexposed areas.
+In localized low-light regions, EAFNet recovers fine structures without introducing noise, enhancing perceptual clarity.
+These visual improvements align with the quantitative margins in Table III, highlighting the advantage of EAFNet within the dual-stream framework.
 Figure 10: Visual comparison of ablation on self-captured videos.
@@
 The luminance histograms show that our method increases the visibility of extremely dark areas without causing an increase in high-luminance pixels.
-V-C Ablation study
-In this section, we investigate different components of our system individually to understand the capability of the system and the source of efficiency.
-Specifically, all quantitative evaluations are performed on Prabhakar’s dataset[30].
+V-D Ablation study
+In this section, we first evaluate the robustness of our framework under different dual-camera deployment conditions using our self-captured real-world dataset, focusing on the impact of calibration accuracy, input ordering, and frame rate.
+We then perform an ablation study on EAFNet to analyze the contribution of each component and the source of efficiency, with all quantitative results reported on Prabhakar’s dataset [44].
+Dual-Camera Analysis.
+In a dual-camera system, imperfect calibration and potential desynchronization can lead to geometric misalignment and motion-induced misalignment, respectively.
+In our prototype, no special effort is made to achieve sub-pixel calibration or perfect synchronization.
+Instead, the fusion framework is designed to tolerate moderate deviations in both geometry and timing.
+The training data include dynamic scenes with motion and occlusion, allowing the model to learn robustness to such spatial inconsistencies.
+Table IV reports t-SSIM under four deployment variants.
+Variations across deployment variants are small for all methods ($\leq$0.003), indicating that, within the tested range, moderate calibration errors, input order changes, and frame-rate reductions have limited impact on temporal flickering.
+Minor fluctuations are likely due in part to the cropping required for alignment rather than alignment failure itself.
+Overall, the results suggest that, within the tested range of deviations, temporal stability can be maintained without the need for precise calibration and strict synchronization, but a higher level of calibration and synchronization may still be beneficial for optimizing performance in more demanding scenarios.
+TABLE IV: t-SSIM comparison under different dual-camera deployment settings.
+base: Baseline dual-camera system settings.
+w/o Cal.: Dual-camera system without calibration.
+Random nr.: Randomized input order of non-reference frames.
+Lower fr.: Lower frame rate for non-reference frame input.
+Method settings Base. w/o Cal. Random nr. Lower fr.
+HDR-Trans(ECCV 2022) [34] 0.8923 0.8922 0.8926 0.8924
+DomainPlus(MM 2022) [62] 0.8991 0.8994 0.8997 0.8985
+SCTNet(ICCV 2023) [46] 0.8868 0.8872 0.8842 0.8871
+SAFNet(ECCV 2024) [27] 0.8851 0.8861 0.8848 0.8836
+EAFNet(Ours) 0.9071 0.9079 0.9072 0.9066
+TABLE V: The ablation study of pre-alignment subnetwork and asymmetric cross-feature fusion subnetwork on Prabhakar’s datasets
+pre align fusion
+Method GLA EFSM ACA $f^{guide}$ $\mu PSNR$-L $\mu SSIM$-L
+1 41.29 39.47 0.9728 0.9888
+2 41.43 39.73 0.9722 0.9890
+3 41.17 39.33 0.9726 0.9780
+4 CA 41.61 39.80 0.9726 0.9892
+5 41.48 39.79 0.9723 0.9893
+6 41.13 39.12 0.9708 0.9887
+7 40.81 38.70 0.9703 0.9784
+Ours 41.80 40.14 0.9731 0.9895
 Global Luminance Alignment.
@@
 The visualization results of the ablation experiments on GLA with Prabhakar’s dataset are shown in Fig.
-13, in the first row, other variants fail to restore the overexposed details of the window frame.
+11, in the first row, other variants fail to restore the overexposed details of the window frame.
 In the second row, ghosting occurs in the overexposed areas outside the window.
@@
 TABLE VI: Ablation investigation about global luminance alignment.
-Channels Original
+Method Channels Original
 Gamma
 Correction
-GLA PSNR-$\mu$ PSNR-$l$ SSIM-$\mu$ SSIM-$l$
-9 $\surd$ $\surd$ $\surd$ 41.80 40.14 0.9731 0.9895
-6 $\surd$ $\surd$ 41.17 39.33 0.9726 0.9780
-6 $\surd$ $\surd$ 41.08 38.97 0.9729 0.9884
-6 $\surd$ $\surd$ 41.19 39.54 0.9729 0.9882
+GLA $\mu PSNR$-$l$ $\mu SSIM$-$l$
+1 6 41.17 39.33 0.9726 0.9780
+2 6 41.08 38.97 0.9729 0.9884
+3 6 41.19 39.54 0.9729 0.9882
+Ours 9 41.80 40.14 0.9731 0.9895
 Exposure-guided Feature Selection Module.
@@
 As shown in Fig.
-12, this integration results in a significant enhancement in detail preservation while the PSNR-$\mu$ is increased by 0.37 dB over the baseline.
+10, this integration results in a significant enhancement in detail preservation while the PSNR-$\mu$ is increased by 0.37 dB over the baseline.
 The luminance histograms show that the joint role of GLA and EFSM increases the visibility of extremely dark areas without causing an increase in high-luminance pixels.
@@
 As shown in Fig.
-14, the proposed GLA and EFSM work collaboratively to fully explore exposure information and enhance meaningful details.
+12, the proposed GLA and EFSM work collaboratively to fully explore exposure information and enhance meaningful details.
 Specifically, the shadow details in low-exposure features are enhanced, while the overexposed regions in high-exposure features are suppressed.
@@
 TABLE VII: Experiments on changing the proportions in exposure-guided feature selection module
-Proportions PSNR-$\mu$ PSNR-$l$ SSIM-$\mu$ SSIM-$l$
-1:2 41.41 39.66 0.97281 0.97907
-1:1 41.43 39.70 0.97284 0.98918
-2:1 41.80 40.14 0.97311 0.98953
-4:1 41.68 40.00 0.97251 0.98934
-8:1 41.72 40.02 0.97255 0.98938
+Method Proportions $\mu PSNR$-$l$ $\mu SSIM$-$l$
+1 1:2 41.41 39.66 0.9728 0.9791
+2 1:1 41.43 39.70 0.9728 0.9892
+3 4:1 41.68 40.00 0.9725 0.9893
+4 8:1 41.72 40.02 0.9726 0.9894
+Ours 2:1 41.80 40.14 0.9731 0.9895
 Asymmetric Cross-feature Fusion Subnetwork.
@@
 14 and Fig.
-15 further confirms that our method effectively mitigates misalignment artifacts.
+13 further confirm that our method effectively mitigates misalignment artifacts.
 Additionally, we evaluate the role of the asymmetric cross-attention (ACA) by comparing it with cross-attention (CA) mechanism.
@@
 Figure 13: Visual comparison of asymmetric cross-attention and cross-attention. Our method reduces misalignment artifacts.
-Restoration Subnetwork In this section, we conduct ablation experiments to evaluate the role of the frequency-domain design and multiscale structure in HDR reconstruction.
+Restoration Subnetwork.
+In this section, we conduct ablation experiments to evaluate the role of the frequency-domain design and multi-scale structure in HDR reconstruction.
 As shown in Table VIII, the baseline model (merging network in AHDNet) achieves a PSNR-$\mu$ of 40.81 and an SSIM-$\mu$ of 0.9719.
 First, we remove the frequency-domain design, and the results show that the PSNR-$\mu$ dropped by 0.09.
-Next, we perform an ablation experiment on the multiscale structure by removing it.
-The results indicate that, without the multiscale structure, the PSNR-$\mu$ decreased by 0.22.
-Our full model, which incorporates both the frequency-domain design and multiscale structure, yields the highest performance, achieving a PSNR-$\mu$ of 41.80 and an SSIM-$\mu$ of 0.97311.
-TABLE VIII: Visual comparison of the separate effects of restoration network.
-model PSNR-µ PSNR-L SSIM-µ SSIM-L
-baseline 40.81 38.60 0.97192 0.98813
-w/o mutil-scale 41.58 39.86 0.97249 0.98911
-w/o DWT 41.71 40.04 0.97286 0.98937
-Ours 41.80 40.14 0.97311 0.98953
+Next, we perform an ablation experiment on the multi-scale structure by removing it.
+The results indicate that, without the multi-scale structure, the PSNR-$\mu$ decreased by 0.22.
+Our full model, which incorporates both the frequency-domain design and multi-scale structure, yields the highest performance, achieving a PSNR-$\mu$ of 41.80 and an SSIM-$\mu$ of 0.97311.
 Figure 14: Visual comparison of ablation on self-captured videos. Our method reduces most ghosting artifacts.
+TABLE VIII: Comparison of the separate effects of the restoration subnetwork.
+Method model $\mu PSNR$-L $\mu SSIM$-L
+1 baseline 40.81 38.60 0.9719 0.9881
+2 w/o multi-scale 41.58 39.86 0.9725 0.9891
+3 w/o DWT 41.71 40.04 0.9729 0.9895
+Ours full model 41.80 40.14 0.9731 0.9895
+Figure 15: Performance variation of different models under simulated parallax.
+The horizontal axis denotes the introduced horizontal disparity in pixels, and the vertical axis shows the performance ($\mathrm{PSNR}$-$\mu$).
+All methods exhibit performance degradation as disparity increases, with SAFNet and DomainPlus showing the most significant drop.
+Our method consistently achieves the highest reconstruction quality across the entire disparity range, demonstrating superior robustness to parallax.
+V-E limitation
+Camera parallax.
+The inherent fixed-baseline nature of our dual-camera system inevitably introduces parallax between the two views, especially for objects at close range.
+Given our hardware configuration ($B=40\mathrm{mm}$, $f=6.5\mathrm{mm}$, $p=3.45\mu\mathrm{m}$, $f_{\mathrm{px}}\approx 1884,\mathrm{px}$), the expected horizontal disparity for a target at distance $Z$ follows
+$\Delta x\approx\frac{f_{\mathrm{px}}B}{Z}.$ (13)
+This corresponds to $\sim 9.4$ pixels for targets at $8\mathrm{m}$ and over $25$ pixels for those at $3\mathrm{m}$.
+Such disparity is inherent to our system geometry and can be mitigated by reducing the camera baseline.
+We analyze the impact of parallax on model performance.
+As shown in Fig.
+15, performance degrades as parallax increases.
+This trend is particularly pronounced for CNN-based approaches (SAFNet and DomainPlus), which are inherently sensitive to spatial misalignments due to their limited receptive fields and lack of explicit geometry modeling.
+Large disparities can cause feature mismatches across views, leading to ghosting, blur, or texture loss in the fused HDR output.
+Nevertheless, our method consistently maintains the highest reconstruction quality across all disparity ranges, demonstrating its robustness in the presence of parallax.
+Efficiency Analysis The overall efficiency of HDR video processing depends on both data acquisition throughput and reconstruction latency.
+Our system addresses the acquisition bottleneck by employing an asynchronous dual-camera design, which decouples the capture of varying-exposure frames from the temporally consistent reference stream.
+This configuration eliminates the delays typically introduced by long high-exposure frames in single-camera systems, thereby enabling real-time HDR frame acquisition.
+Although the proposed reconstruction model (EAFNet) is not explicitly optimized for real-time inference, it demonstrates competitive runtime performance in practice, as summarized in Table III.
+The system design thus provides a practical balance between capture efficiency and reconstruction quality, supporting high-rate video acquisition with high dynamic range and temporal consistency.
 VI CONCLUSION
-In this work, we propose a dual-camera system (DCS) to mitigate the temporal inconsistency with AE methods, and existing MEF methods can be seamlessly extended to our system with temporally consistent output.
-In addition, we introduce an exposure adaptive fusion network with two novel subnetworks to provide a more stable and efficient solution for video HDR reconstruction.
-Experimental results demonstrate that our approach outperforms existing methods, offering stable performance and high-quality reconstruction in complex real-world video environments.
-Our dual-camera framework provides a promising direction for real-time HDR video capture.
+In this work, we revisit the fundamental cause of temporal instability in alternating-exposure (AE) HDR video, which lies in the entanglement of temporal luminance anchoring with exposure-dependent detail selection, and propose a dual-stream paradigm that explicitly decouples these two roles.
+Specifically, we introduce an exposure-asymmetric dual-camera system (DCS) that fixes the reference exposure to stabilize frame-to-frame luminance while allowing the auxiliary stream to vary exposure for capturing high dynamic range details.
+To support this system, we design EAFNet to effectively suppress ghosting artifacts while preserving fine details in real-world scenes.
+Extensive experiments on multiple datasets and real-world sequences demonstrate that the proposed paradigm improves both temporal stability and reconstruction quality compared with AE-based baselines and HDR image deghosting methods, while remaining cost-efficient and deployment-friendly.
+Our dual-stream solution provides a promising direction for real-time HDR video capture.